### v3

Abstract
We consider separable nonconvex optimization problems under affine constraints. For these problems, the Shapley-Folkman theorem provides an upper bound on the duality gap as a function of the nonconvexity of the objective functions, but does not provide a systematic way to construct primal solutions satisfying that bound. In this work, we develop a two-stage approach to do so. The first stage approximates the optimal dual value with a large set of primal feasible solutions. In the second stage, this set is trimmed down to a primal solution by computing (approximate) Carathéodory representations. The main computational requirement of our method is tractability of the Fenchel conjugates of the component functions and their (sub)gradients. When the function domains are convex, the method recovers the classical duality gap bounds obtained via Shapley-Folkman. When the function domains are nonconvex, the method also recovers classical duality gap bounds from the literature, based on a more general notion of nonconvexity.
Keywords:
Shapley-Folkman Carathéodory nonconvex optimization duality Frank-Wolfe
MSC:
90C26 90C46 65K05
1.   Introduction
We consider nonconvex separable optimization problems of the form
$\begin{split}\mbox{minimize}&\ \sum_{i=1}^{n}f_{i}(x_{i})\\ \mbox{subject to }&\ \sum_{i=1}^{n}A_{i}x_{i}\leq b,\\ &x_{i}\in X_{i},\ i=1,\dots,n,\end{split}$ (P)
in the variables $x_{i}\in\mathbb{R}^{d_{i}}$, with $A_{i}\in\mathbb{R}^{m\times d_{i}}$, $b\in\mathbb{R}^{m}$. We make the following assumption on the structure of the functions considered.
Assumption 1.1
For all $i\in\{1,\dots,n\}$, $f_{i}$ is proper, lower semicontinuous and has an affine minorant. Moreover $X_{i}=\operatorname{dom}f_{i}$ is compact.
Note that we do not assume convexity of either the functions $f_{i}$ or their domains $X_{i}$. We also assume that one can efficiently compute the conjugates of the functions $f_{i}$ and their gradients (or subgradients). We let $d=\sum_{i=1}^{n}d_{i}$ and $A=[A_{1}\dots A_{n}]\in\mathbb{R}^{m\times d}$ so that the constraints can be expressed in the more compact form $Ax\leq b$, where $x=[x_{1}^{\top}x_{2}^{\top}\dots x_{n}^{\top}]^{\top}$.
1.1.   Motivation
Real-world problems with the structure of (P) arise, for example, in power system management bertsekas1983optimal ; vujanic2016decomposition , supply chain optimization vujanic2014large , resource allocation udell2016bounding , network utility problems bi2016refined and feature selection askari2024naive . Let us illustrate one such popular application, commonly referred to as the unit commitment problem bertsekas1983optimal . In this problem, an operator manages a network of $n$ plants producing electricity, and must meet a certain electricity demand (given ahead of time) over $m$ time steps, say $D_{1},\dots,D_{m}$. At each time step $t$, the operator can choose how much power $g_{i}^{t}$ is produced by a plant $i$. The constraint on the demand can then be written as
$\sum_{i=1}^{n}g_{i}^{t}\geq D_{t},\quad t=1,\dots,m.$ (1)
For each plant $i$, the production profile over $t$ time steps $g_{i}=[g_{i}^{1},\dots,g_{i}^{t}]$ has a certain associated cost $c_{i}(g_{i})$. The goal of the operator is then to minimize the overall cost, i.e.
$\mbox{minimize}\ \sum_{i=1}^{n}c_{i}(g_{i}).$ (2)
Finally, at each time step, a plant $i$ is either turned off, or turned on and producing between a minimum and a maximum capacity. This makes the domain of the objective functions non-convex. We have thus described a non-convex problem whose objective function (2) is separable and whose constraints (1) are affine, just as in (P). We will come back in more details to the unit commitment problem in the numerical section of the paper.
1.2.   Background
The Shapley-Folkman theorem states that the Minkowski sum of arbitrary sets gets increasingly close to its convex hull as the number of terms in the sum increases. This was first established by Shapley and Folkman in private communications, and first published in starr1969quasi . This result can be used to show that the duality gap of (P) vanishes when the number $n$ of component functions goes to infinity, while the dimension of the constraints $m$ remains bounded aubin1976estimates . More precisely, the duality gap can be bounded by $m\max_{i}\rho(f_{i})$, where $\rho(f_{i})$ is a measure of the nonconvexity of the function $f_{i}$ aubin1976estimates ; bi2016refined ; udell2016bounding . In cases where the domain of the component functions is not convex, the Shapley-Folkman theorem can also be used to prove similar bounds on the duality gap, this time based on more general measures of nonconvexity than the ones mentioned above bertsekas2014constrained ; vujanic2016decomposition . Those duality gap bounds are particularly interesting in many real-world applications where $n>>m$. This is often the case for the unit commitment problem described in Section 1.1 where $n$, the number of plants producing electricity, is typically much larger than $m$, the number of time steps at which the demand has to be satisfied.
1.3.   Roadmap of the paper and contributions
The goal of Section 2 is to provide the reader with all the necessary mathematical tools to present duality gap bounds for problems of the form (P). We emphasize that those bounds are well-known, see for example aubin1976estimates ; ekeland1999convex , and are not part of our contributions. They are however central for the main contribution of the paper. These bounds are based on the Shapley-Folkman theorem, which is itself based on the Carathéodory decomposition results. After introducing some important definitions, we therefore recall the Carathéodory results, in both the convex and conic cases.
Our first contribution comes at this stage where we adapt the results of wirth2024frank to provide and analyze an algorithm which computes the conic Carathéodory representation of a given vector. These results allow us to recall a proof of the Shapley-Folkman theorem. We end Section 2 by presenting different (but similar in flavor) duality gap bounds for (P), notably the central result from Aubin and Ekeland aubin1976estimates .
The bounds on the duality gap presented in Section 2 are not constructive, in the sense that the proofs do not provide a solution meeting those bounds. The main contribution of this work is a systematic approach to obtain solutions of (P) meeting the Shapley Folkman duality gap bounds. We detail this approach in Section 3, and show that it can be broken down into two main stages.
The first stage, described in section 3.1, is a conditional gradient method on a smooth convex objective aiming at approximating the optimal dual value. This method converges sublinearly and, crucially here, we show that the linear minimization oracle comes down to computing the Fenchel conjugate of the functions $f_{i}$, a fairly routine step. The result of the algorithm can be cast as a conic combination of extreme points of the primal feasible set.
The second stage of the method, described in Section 3.2, requires trimming this conic combination to a conic Carathéodory representation using only $n+m+1$ elements. This can be done using the algorithm derived in Section 2, justifying our first contribution. When the domains of the functions $X_{i}$ are convex, we show in Section 3.3 that this directly yields a primal feasible solution meeting the classical bounds of aubin1976estimates ; ekeland1999convex presented in Section 2, up to an arbitrarily small additive term. This is summarized in Theorem 3.2, which we consider to be the main result of this work. In the case of nonconvex domains, we show that the obtained solution is not feasible on at most $m+1$ of the $n$ component functions. Under various additional assumptions, we show that a final reconstruction step can be carried out to obtain fully feasible solutions, and we derive duality gap bounds similar to the ones above, this time based on a more general notion of nonconvexity bertsekas2014constrained ; vujanic2014large ; vujanic2016decomposition . The main results for nonconvex domains are stated in Theorem 3.3.2.2 and Theorem 3.1.
Let us emphasize that the main requirements of the method described in Section 3 are (i) the computation of the optimal dual value of (P) (ii) being able to compute Fenchel conjugates of the functions $f_{i}$. These fairly routine computational assumptions make our method versatile and applicable in many real-world settings.
In Section 4, we discuss a variant of our method described in Section 3 to solve problem (P). As mentioned earlier, the second stage of our method requires computing a Carathéodory representation of the output of the first stage, and this can be done using the algorithm presented in Section 2. However, this algorithm has high complexity and can therefore be quite slow on large-scale instances of (P). Inspired by combettes2023revisiting , we show in Section 4 that one can use conditional gradient methods to obtain an approximate Carathéodory representation of the output of the first stage instead of an exact one. Those methods have a fast linear rate of convergence. Moreover, we show that using such a conditional gradient method also reduces the dependency on $m$ in the duality gap bounds of the reconstructed solutions, at the price of an exponentially decreasing error term, which depends on the number of iterations of the conditional gradient method.
Finally, in Section 5, we implement and test our method on two numerical applications, comparing final performances and running times of computing exact versus approximate Carathéodory representations.
1.4.   Related work
When the underlying domain of the functions are convex, udell2016bounding describes a randomized algorithm achieving the bounds derived from Shapley-Folkman with probability one. To the best of our knowledge, this is the only other work describing a method to build solutions matching these bounds. However, the approach requires knowledge of the full solution space of the convex relaxation of (P), as opposed to our method which only requires access to the optimal dual value. We give a further comparison of our method with the method of udell2016bounding in Section 2.4, where the tools necessary for a further discussion are introduced.
In the special case of mixed-integer linear programs (MILPs) with unique solution, vujanic2014large ; vujanic2016decomposition describe algorithms achieving the bounds for nonconvex domains derived from Shapley-Folkman under various additional assumptions on the domain.
Let us also mention here that the bounds derived from Shapley-Folkman can be slightly improved. In bi2016refined , the authors prove duality gap bounds based on a slightly more refined version of the nonconvexity of a function, known as $k$-th nonconvexity. Although our results immediately extend to this more refined version, we omit them to simplify the exposition.
In kerdreux2017approximate , the authors describe an approach to obtain approximate
Carathéodory representations with high sampling ratios, which they use to prove approximate Shapley-Folkman theorems. Those, in turn, can be used to prove duality gap bounds which reduce the dependence on the constraint dimension $m$, at the price of an additional term due to the approximation error. The approach is, however, not constructive, and the duality gap bounds provided are on a perturbed version of the primal problem.
Finally, let us briefly mention branch and bound algorithms. While a complete overview is beyond the scope of this work (we refer the reader to lawler1966branch ; horst2013global ; tawarmalani2013convexification for a thorough exposition), these algorithms form a widely used family of methods for tackling nonconvex problems. We therefore highlight the main differences between these methods and our approach. Branch and bound algorithms aim to find exact solutions of the problem by recursively partitioning the feasible set and computing upper and lower bounds on the optimal function value within each partition. In particular, the design of the partitioning step is a key aspect of these methods, and must often be tailored to the problem at hand. Moreover, computing the bounds also depends on the structure of the problem, and the number of partitions required to obtain a solution can be prohibitively large in high dimension. On the other hand, our method is aimed at finding good approximate solutions in reasonable time, and relies on fairly general assumptions (computation of Fenchel conjugates), making it applicable without further specification in a variety of settings.
2.   Preliminaries
For $k\in\mathbb{N}$, $\Delta_{k}$ is the simplex of dimension $k-1$. For a set $S$, we write its closure $\operatorname{cl}(S)$, its convex hull $\operatorname{conv}(S)$ and the set of its extreme points $\operatorname{ext}(S)$. For a vector $x\in\mathbb{R}^{d}$, we define $\left\lVert x\right\rVert_{+}=\sqrt{\sum_{j=1}^{d}\max\left(x_{j},0\right)^{2}}$.
If $f$ is a function not everywhere +$\infty$+ ∞, with an affine lower bound, the (Fenchel) conjugate of $f$ is defined as
$f^{*}(y)=\sup_{x\in\operatorname{dom}f}\{y^{\top}x-f(x)\}.$ (3)
The biconjugate of $f$, defined as the conjugate of $f^{*}$ and written $f^{**}$, is also the pointwise supremum of all affine functions minorizing $f$ (hiriart1996convex, , Theorem X.1.3.5). In particular, we have $\operatorname{epi}f^{**}=\operatorname{cl}\operatorname{conv}\operatorname{epi}f$. Moreover, when $f$ is lower semicontinuous and 1-coercive, $\operatorname{conv}\operatorname{epi}f$ is closed and hence $\operatorname{epi}f^{**}=\operatorname{conv}\operatorname{epi}f$ (hiriart1996convex, , Theorem X.1.5.3).
The nonconvexity of a function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined as aubin1976estimates
$\rho(f)=\sup\left\{f\left(\sum_{j=1}^{p}\alpha_{j}x_{j}\right)-\sum_{j=1}^{p}\alpha_{j}f(x_{j})\mid p\in\mathbb{N},x_{j}\in\operatorname{dom}f,\alpha\in\Delta_{p}\right\}.$ (4)
Clearly $\rho(f)=0$ if $f$ is convex. Note also that in the above definition, we do not require $\sum_{j=1}^{p}\alpha_{j}x_{j}$ to be in $\operatorname{dom}f$. In particular we then have $\rho(f)=\infty$ if $\operatorname{dom}f$ is not convex. Under Assumption 1.1, the nonconvexity $\rho(f)$ can also be expressed in terms of the biconjugate of $f$ (bi2016refined, , Proposition 6),
$\rho(f)=\sup_{x\in\operatorname{dom}f^{**}}\{f(x)-f^{**}(x)\}.$
2.1.   Dual and convex relaxation
Following kerdreux2017approximate we define, for $\lambda\in\mathbb{R}^{m}$,
$\Psi(\lambda)=\inf_{x_{i}\in X_{i}}\left\{\sum_{i=1}^{n}f_{i}(x_{i})+\lambda^{\top}(\sum_{i=1}^{n}A_{i}x_{i}-b)\right\}$ (5)
The Lagrange dual of (P) is given by
$\begin{split}\mbox{maximize}&\ \Psi(\lambda)\\ \text{subject to }&\ \lambda\geq 0.\end{split}$ (D)
The bidual of (P), which is the dual of (D), is then given by
$\begin{split}\mbox{minimize}&\ \sum_{i=1}^{n}f_{i}^{**}(x_{i})\\ \mbox{subject to }&\ Ax\leq b,\\ &\ x_{i}\in\operatorname{conv}(X_{i}),\quad i=1,\dots,n.\end{split}$ (CoP)
The fact that the domain is now $\operatorname{conv}(X_{i})$ follows from $\operatorname{dom}f_{i}^{**}=\operatorname{conv}(\operatorname{dom}f_{i})$. From weak duality, the optimal value of (D) is a lower bound on the optimal value of (P). The following proposition from lemarechal2001geometric gives a sufficient condition for the optimal values of (D) and (CoP) to match.
Proposition 2.1
Assume $\Psi$ is not everywhere $\infty$- ∞ and there exists a feasible $\hat{x}$ in the interior of $\operatorname{dom}\sum_{i=1}^{n}f_{i}^{**}$. Then strong duality holds, i.e. the optimal values of (D) and (CoP) match.
From now on, we assume that strong duality holds and let
$v^{*}:=\max_{\lambda\geq 0}\Psi(\lambda).$ (6)
In the rest of the paper, we assume that one can efficiently obtain $v^{*}$, by either solving (D) or (CoP). This is a relatively fair assumption, as both (D) and (CoP) are convex problems, for which standard convex optimization solvers can be used.
2.2.   Carathéodory
Carathéodory’s theorem and its conic version are core ingredients in proving the Shapley-Folkman theorem and the resulting duality gap bounds. We recall both results below, and also present the proof of the conic version, as it will be useful later in the paper. The classical Carathéodory theorem states that a vector in the convex hull of a set in $\mathbb{R}^{p}$ can be written as a convex combination of at most $p+1$ points of the original set.
Theorem 2.1 (Carathéodory)
Let $S\subset\mathbb{R}^{p}$ and $w\in\operatorname{conv}S$. Then there exists $\{\lambda_{j}\}_{j=1}^{p+1}\subset\mathbb{R}_{+}$, $\{w_{j}\}_{j=1}^{p+1}\subset S$ with $\sum_{j=1}^{p+1}\lambda_{j}=1$, such that $w=\sum_{j=1}^{p+1}\lambda_{j}w_{j}$.
Proof
See (rockafellar2015convex, , Theorem 17.1).
Let $\operatorname{Po}(S)=\left\{\sum_{j}\lambda_{j}v_{j}\mid\lambda_{j}\geq 0,v_{j}\in S\right\}$ be the conic hull of $S$. The conic Carathéodory theorem states that any point in the conic hull can be written as a conic combination of at most $p$ points in the original set.
{restatable}
[Conic Carathéodory]thmConicCaratheodory Let $S\subset\mathbb{R}^{p}$ and $w\in\operatorname{Po}S$. Then there exists $\{\lambda_{j}\}_{j=1}^{p}\subset\mathbb{R}_{+}$, $\{w_{j}\}_{j=1}^{p}\subset S$, such that $w=\sum_{j=1}^{p}\lambda_{j}w_{j}$.
Proof
Since $w\in\operatorname{Po}(S)$, there exists $N\in\mathbb{N}$, $\lambda\in\mathbb{R}^{N}_{+}$ and $w_{1},\dots,w_{N}\in S$ such that $w=\sum_{j=1}^{N}\lambda_{j}w_{j}$. We assume $N>p$ (if $N\leq p$, we are already done). We claim that we can rearrange the coefficients so that one of the $w_{j}$’s has coefficient 00. To do so, suppose $\lambda_{j}>0$ for all $j=1,\dots,N$ (otherwise we are done). Since $N>p$, the $w_{j}$’s are linearly dependent, so there exists $\delta_{1},\dots,\delta_{N}$ not all zero such that
$\sum_{j=1}^{N}\delta_{j}w_{j}=0.$
Let $t^{*}=\max\{t\geq 0\mid\lambda_{j}-t\delta_{j}\geq 0\text{ for all $j=1,\dots,N$}\}$ and set $\lambda_{j}^{\prime}=\lambda_{j}-t^{*}\delta_{j}$ for all $j$. Then $\lambda_{j}^{\prime}\geq 0$ for all $j$. Moreover, there exists $j_{0}$ such that $\lambda_{j_{0}}^{\prime}=0$ and
$\sum_{j=1}^{N}\lambda_{j}^{\prime}w_{j}$ $=\sum_{j=1}^{N}\lambda_{j}w_{j}-t^{*}\sum_{j=1}^{N}\delta_{j}w_{j}=w,$
which proves that we can indeed reduce the number of strictly positive coefficients to $N-1$. This argument can be repeated for $r$ iterations as long as $N-r>p$, which proves the theorem.
The remainder of this subsection is dedicated to the design of an algorithm for computing conic Carathéodory representations. We consider a vector $w^{*}\in\mathbb{R}^{p}$, and a subset $S\subset\mathbb{R}^{p}$ of size $N$ which we write $S=\{w_{1},\dots,w_{N}\}$, and we assume that $w^{*}\in\operatorname{Po}(S)$. In other words, for coefficients $\lambda_{1},\dots,\lambda_{N}>0$, we have
$w^{*}=\sum_{j=1}^{N}\lambda_{j}w_{j}.$ (7)
We assume that the vectors $w_{1},\dots,w_{N}$ and the coefficients $\lambda_{1},\dots,\lambda_{N}$ are known. Let us give some intuition behind the algorithm to reduce the number of components from $N$ to $p$, which is also the main idea behind the proof of Theorem 2.2. Since $N>p$, the vectors $w_{1},\dots,w_{N}$ are linearly dependent. Finding the coefficients for which the resulting linear combination is 0 entails solving a linear system in dimension $\mathbb{R}^{p\times N}$. Since this linear system has a trivial solution (namely, setting all coefficients to 0), an additional random row ensures that non-trivial solutions are obtained (with probability one).
Once the non-trivial solution $\delta$ has been found, a simple algebraic trick shows that one can subtract the (scaled) solution $\delta$ from the original $\lambda$ conic coefficients, and that the resulting conic coefficients have at least one coordinate equal to 0, which we can thus remove from the conic representation. Iterating this procedure will yield a conic representation with precisely $p$ elements. Now, instead of considering all of the vectors $w_{1},\dots,w_{N}$ in the first step above, it is enough to only consider $p+1$ of them, as those are also linearly dependent, and this allows to reduce the dimension of the linear system to be solved at each iteration to $\mathbb{R}^{(p+1)\times(p+1)}$. We summarize the method in Algorithm 1. For a matrix $W$, we write $W_{:,k}$ as its $k$-th column and $W_{k,:}$ as its $k$-th row. Moreover, $W_{:,k_{1}:k_{2}}$ is the submatrix of $W$ composed of the $k_{2}-k_{1}+1$ columns between $k_{1}$ and $k_{2}$.
Algorithm 1 Exact Carathéodory
1:  Input: $w_{1},\dots,w_{N}$ and nonnegative coefficients $\lambda_{1},\dots,\lambda_{N}$ such that $w^{*}=\sum_{j=1}^{N}\lambda_{j}w_{j}$.
2:  Build $W\in\mathbb{R}^{p\times N}$, the matrix whose columns are the vectors $w_{j}$.
3:  Draw vector $u\in\mathbb{R}^{N}$ with uniform distribution on the unit sphere and let $\tilde{W}_{u}=\begin{pmatrix}W\\ u^{\top}\end{pmatrix}$.
4:  $\tilde{W}^{s}_{u}\leftarrow(\tilde{W}_{u})_{:,1:p+1}$
5:  $\alpha\leftarrow\lambda_{1:p+1}$
6:  $c\leftarrow p+2$
7:  while $c\leq N+1$ do
8:     Solve $(\tilde{W}^{s}_{u})\delta=\begin{pmatrix}\mathbf{0}_{p}\\ 1\end{pmatrix}$
9:     $t^{*}\leftarrow\min\{\frac{\alpha_{j}}{\delta_{j}}\mid\delta_{j}>0\}$
10:     $\alpha^{\prime}\leftarrow\alpha-t^{*}\delta$
11:     $j^{*}\leftarrow\operatorname*{arg\,min}_{j}\alpha_{j}^{\prime}$
12:     Remove element at index $j^{*}$ from $\alpha^{\prime}$ and column $j^{*}$ from $\tilde{W}^{s}_{u}$
13:     if $c=N+1$ then
14:        $\alpha\leftarrow\alpha^{\prime}$
15:        Break
16:     else
17:        $\alpha\leftarrow(\alpha^{\prime},\lambda_{c})$
18:        $\tilde{W}^{s}_{u}\leftarrow\left(\tilde{W}^{s}_{u},(\tilde{W}_{u})_{:,c}\right)$
19:        $c\leftarrow c+1$
20:     end if
21:  end while
22:  Return $\alpha,(\tilde{W}^{s}_{u})_{1:p,:}$.
Theorem 2.2
With probability 1, Algorithm 1 outputs $\alpha\in\mathbb{R}^{p}_{+}$ and $p$ vectors $\{w_{j_{l}}\}_{l=1}^{p}\subset S$, such that
$w^{*}=\sum_{l=1}^{p}\alpha_{l}w_{j_{l}}.$ (8)
Proof
Our proof basically mirrors the one of Theorem 2.2. By the definition of $W$ in Algorithm 1, we have $W\lambda=w^{*}$. Let $W^{s}$ be the matrix $\tilde{W}^{s}_{u}$ from which we have removed the last row (namely the row composed of elements of the random vector $u$). At the beginning of the first iteration, we then have
$\ W\lambda=W^{s}\alpha+(W_{:,c:N})\lambda_{c:N},$
$\Rightarrow$ $\ W^{s}\alpha=w^{*}-(W_{:,c:N})\lambda_{c:N}.$ (9)
We claim that (9) is true at the beginning of every iteration. We saw that it was true at the beginning of the first iteration, now suppose it is true at the beginning of some iteration.
Since $W^{s}\in\mathbb{R}^{p\times(p+1)}$, the system $W^{s}\delta=\mathbf{0}_{p}$ must have a non-trivial solution. Let $\delta^{*}$ be one such solution. Observe that
$\tilde{W}^{s}_{u}\frac{\delta^{*}}{((\tilde{W}^{s}_{u})_{p+1,:})^{\top}\delta^{*}}=\begin{pmatrix}\mathbf{0}_{p}\\ 1\end{pmatrix},$
and hence a solution of the system on line 8 exists as long as $((\tilde{W}^{s}_{u})_{p+1,:})^{\top}\delta^{*}\not=0$. Since the last row $\tilde{W}^{s}_{u}$ has uniform distribution on the unit sphere, this holds with probability 1.
One can then check that lines 9-11 in the algorithm are simple computations to find $\alpha^{\prime}\geq 0$ such that $W^{s}\alpha^{\prime}=W^{s}\alpha$ and $\alpha_{j^{*}}^{\prime}=0$. Removing the index $j^{*}$ from $\alpha^{\prime}$ and the column $j^{*}$ from $W^{s}$, we get $\alpha^{\prime}\in\mathbb{R}^{p}$, $W^{s}\in\mathbb{R}^{p\times p}$ and
$W^{s}\alpha^{\prime}=w^{*}-(W_{:,c:N})\lambda_{c:N}.$
At this point two cases arise.
Case 1: If $c=N+1$, it must be that $W^{s}\alpha^{\prime}=w^{*}$, and we have therefore found a conic combination of $w^{*}$ with $p$ coefficients, so we are done.
Case 2: If $c<N+1$, we can append $\lambda_{c}$ to $\alpha^{\prime}$ and column $c$ of $\tilde{W}$ to $\tilde{W}^{s}$, as on lines 17 and 18. This results in $\alpha\in\mathbb{R}^{p+1}$ and $W^{s}\in\mathbb{R}^{p\times(p+1)}$ such that
$W^{s}\alpha=w^{*}-(W_{:,c+1:N})\lambda_{c+1:N}.$
Updating $c\leftarrow c+1$ as on line 19, and starting a new iteration, we are back to the equation (9), except that we have reduced the column dimension of the matrix $W_{:,c:N}$ by $1$. We can iterate this argument until case 1 above is true.
Note that similar algorithms to obtain Carathéodory representations have been developed in beck2017linearly and in wirth2024frank . Those approaches are however aimed at reducing the size of the active set of conditional gradient methods, and hence are looking for convex representations via the classic Carathéodory result (Theorem 2.1). In contrast, we are looking for representations in the conic hull of $S$ via the proof of Theorem 2.2. Our approach is closer to the one of wirth2024frank , in the sense that we make no assumption on how the linear system is solved on line 8. In particular, when the vectors in $S$ are sparse, efficient direct or indirect sparse solvers may be used. In contrast, the approach in beck2017linearly requires maintaining a matrix in row-echelon form which, depending on the application, might not be computationally efficient. Without further assumption on the structure of the linear systems, we suggest a strategy to achieve complexity $O(p^{2})$ at each iteration, which is similar to the complexity achieved in beck2017linearly and in wirth2024frank .
First, note that at each iteration, the linear system to be solved is the same as the one at the previous iteration, up to one column deletion and one column insertion. One can therefore compute the QR decomposition of $\tilde{W}^{s}_{u}$ on line 4, solve the linear systems on line 8 using the decomposition, and update the QR decomposition by either deleting a column (line 12) or by appending a column (line 18). We give the complexity of this strategy in the next lemma.
Lemma 2.1
Using the above strategy, the complexity of Algorithm 1 is $O(Np^{2})$.
Proof
The complexity of computing the first QR decomposition is $O(p^{3})$. The complexity of solving the linear system given the QR decomposition is $O(p^{2})$, while the complexity of deleting/appending a column is also $O(p^{2})$ (golub2013matrix, , Section 6.5.2). We must do the above two tasks at each iteration, and there are $N-p$ iterations in total. Since $N\geq p$, the complexity is indeed $O(Np^{2})$.
2.3.   The Shapley Folkman theorem
Both versions of Carathéodory’s theorem are central to the proof of the Shapley Folkman theorem, which states that the Minkowski sum of arbitrary sets gets increasingly closer to its convex hull as the number of terms grows. Established in private communications between Shapley and Folkman, it was first published in starr1969quasi .
Theorem 2.3 (Shapley Folkman)
Let $V_{i}\subset\mathbb{R}^{d}$, $i=1,\dots,n$ and
$x\in\sum_{i=1}^{n}\operatorname{conv}(V_{i})$. Then for some set $S\subset\{1,\dots,n\}$ with $|S|\leq d$, we have
$x\in\sum_{i\in\{1,\dots,n\}\setminus S}V_{i}+\sum_{i\in S}\operatorname{conv}(V_{i})$
Proof
Since $x\in\sum_{i=1}^{n}\operatorname{conv}(V_{i})$, by classic Carathédory we have
$x=\sum_{i=1}^{n}\sum_{j=1}^{d+1}\lambda_{ij}v_{ij}$
with $v_{ij}\in V_{i}$ for all $i$, and $\sum_{j=1}^{d+1}\lambda_{ij}=1$ for all $i$. We can then write
$\begin{pmatrix}x\\ \mathbf{1}_{n}\end{pmatrix}$ $=\sum_{i=1}^{n}\sum_{j=1}^{d+1}\lambda_{ij}\begin{pmatrix}v_{ij}\\ e_{i}\end{pmatrix}.$
This is a conic combination in $\mathbb{R}^{n+d}$, so by conic Carathéodory, there exists $\mu\in\mathbb{R}^{n(d+1)}_{+}$ with at most $n+d$ nonzero such that
$\begin{pmatrix}x\\ \mathbf{1}_{n}\end{pmatrix}$ $=\sum_{i=1}^{n}\sum_{j=1}^{d+1}\mu_{ij}\begin{pmatrix}v_{ij}\\ e_{i}\end{pmatrix}.$
In particular, for each $i$, there must be at least one of the $\mu_{ij}$ not equal to 00. Since there are only $n+d$ nonzero coefficients, this implies that for at least $n-d$ indices, the convex combination is trivial, i.e. has only one nonzero coefficient, which is equal to one. For the remaining $d$ indices, the convex combination is non-trivial. In other words, for a set $S$ with $|S|\leq d$,
$x\in\sum_{i\not\in S}V_{i}+\sum_{i\in S}\operatorname{conv}(V_{i}).$
2.4.   Duality gap estimation
The Shapley-Folkman theorem established above is one of the main ingredients for proving duality gap bounds for problem (P). Let us first state an important lemma, which will be useful throughout this section.
Lemma 2.2
Let $v^{*}$ be the dual optimal value as defined in (6). There exists a set $S\subset\{1,\dots,n\}$ of cardinality at most $m+1$ and a vector $\bar{x}\in\mathbb{R}^{d}$, such that
$\bar{x}_{i}\in X_{i}\ \text{ for }i\not\in S,$
$\bar{x}_{i}\in\operatorname{conv}X_{i}\ \text{ for }i\in S,$
$A\bar{x}\leq b,$
and
$v^{*}=\sum_{i\in S}f_{i}^{**}(\bar{x}_{i})+\sum_{i\not\in S}f_{i}(\bar{x}_{i})$
Proof
Since $v^{*}$ is the optimal value of (CoP), there exists $\tilde{x}\in\mathbb{R}^{d}$ with $\tilde{x}_{i}\in\operatorname{conv}X_{i}$ for all $i=1,\dots n$, $A\tilde{x}\leq b$, and $v^{*}=\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})$. Let
$\tilde{z}=\begin{pmatrix}\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})\\ b\end{pmatrix}$
we have
$\tilde{z}$ $\in\sum_{i=1}^{n}\left\{\begin{pmatrix}f_{i}^{**}(x_{i})\\ A_{i}x_{i}\end{pmatrix}\mid x_{i}\in\operatorname{conv}X_{i}\right\}+\mathbb{R}^{m+1}_{+}$
[…]
We can now apply the Shapley-Folkman theorem to show that there exists a set $S$ of cardinality at most $m+1$ such that
$\tilde{z}$ $\in\sum_{i\not\in S}\left\{\begin{pmatrix}f_{i}(x_{i})\\ A_{i}x_{i}\end{pmatrix}\mid x_{i}\in X_{i}\right\}+\sum_{i\in S}\operatorname{conv}\left\{\begin{pmatrix}f_{i}(x_{i})\\ A_{i}x_{i}\end{pmatrix}\mid x_{i}\in X_{i}\right\}+\mathbb{R}^{m+1}_{+}$
$=\sum_{i\not\in S}\left\{\begin{pmatrix}f_{i}(x_{i})\\ A_{i}x_{i}\end{pmatrix}\mid x_{i}\in X_{i}\right\}+\sum_{i\in S}\left\{\begin{pmatrix}f_{i}^{**}(x_{i})\\ A_{i}x_{i}\end{pmatrix}\mid x_{i}\in\operatorname{conv}X_{i}\right\}+\mathbb{R}^{m+1}_{+},$
where we applied Lemma A.2 again. In other words, there exists $\bar{x}\in\mathbb{R}^{d}$, $\bar{x}_{i}\in X_{i}$ for $i\not\in S$ and $\bar{x}_{i}\in\operatorname{conv}X_{i}$ for $i\in S$, and $\delta_{1}\geq 0,\delta_{2}\in\mathbb{R}^{m}_{+}$ such that
$v^{*}=\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})$ $=\sum_{i\not\in S}f_{i}(\bar{x}_{i})+\sum_{i\in S}f_{i}^{**}(\bar{x}_{i})+\delta_{1},$
$b$ $=A\bar{x}+\delta_{2}.$
In particular, $A\bar{x}\leq b$, and since $f_{i}^{**}\leq f_{i}$, we must actually have $\delta_{1}=0$ and
$\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})=\sum_{i\not\in S}f_{i}(\bar{x}_{i})+\sum_{i\in S}f_{i}^{**}(\bar{x}_{i}).$
2.4.1.   Convex domain, nonconvex objectives
When dealing with convex domains $X_{i}$, we have $X_{i}=\operatorname{conv}(X_{i})$. We now review some results covering duality gap estimation. We first recall the original result of aubin1976estimates (see also ekeland1999convex ), giving a bound on the duality gap which depends on the dimension of the constraints $m$ and the nonconvexity of the considered functions, but which is independent of the number of functions $n$.
{restatable}
thmNonConvexDualityGap Let $v^{*}$ be the dual optimal value defined in (6) and $\rho(f_{i})$ be as defined in (4). Assume that $X_{i}$ is convex for all $i$. There exists $x^{*}\in\mathbb{R}^{d}$ such that $Ax^{*}\leq b$ and
$v^{*}\leq\sum_{i=1}^{n}f_{i}(x_{i}^{*})\leq v^{*}+(m+1)\max_{i}\rho(f_{i}).$
Proof
By Lemma 2.2, there exists $\bar{x}\in\mathbb{R}^{d}$ and set $S\subset\{1,\dots,n\}$ of cardinality at most $m+1$ such that $A\bar{x}\leq b$, $\bar{x}_{i}\in X_{i}$ for $i\not\in S$ and $\bar{x}_{i}\in\operatorname{conv}X_{i}=X_{i}$ (since $X_{i}$ is convex by assumption) such that
$v^{*}=\sum_{i\not\in S}f_{i}(\bar{x}_{i})+\sum_{i\in S}f_{i}^{**}(\bar{x}_{i})$
Setting $x^{*}=\bar{x}$, observe that $x^{*}$ is primal feasible. Moreover,
$v^{*}$ $\leq\sum_{i=1}^{n}f_{i}(x_{i}^{*})$
$=\sum_{i\not\in S}f_{i}(x_{i}^{*})+\sum_{i\in S}f_{i}(x_{i}^{*})$
$=v^{*}+\sum_{i\in S}(f_{i}(x_{i}^{*})-f_{i}^{**}(x_{i}^{*}))$
$\leq v^{*}+(m+1)\max_{i}\rho(f_{i}).$
The above theorem is an existence result, and the proof does not provide a way to construct a solution satisfying the bound.
In udell2016bounding , the authors provide an algorithm achieving the bound of Theorem 2.4.1 with probability 1, by solving a randomized version of the convex relaxation (CoP). Although they claim otherwise, their analysis implicitly assumes that the function domains $X_{i}$ are convex. Indeed, their randomized approach yields a solution $x_{i}\in\operatorname{conv}X_{i}$, $i=1,\dots,n$. On a certain index set $\mathcal{J}$ of cardinality at least $n-m$, they show that $i\in\mathcal{J}$ implies that $x_{i}$ is an extreme point of $\operatorname{conv}X_{i}$. This in turn implies that $x_{i}\in X_{i}$ and $f_{i}(x_{i})=f_{i}^{**}(x_{i})$. On non-extreme points, i.e. for $i\not\in\mathcal{J}$, the authors bound $f_{i}(x_{i})-f_{i}^{**}(x_{i})$ by $\rho(f_{i})$. However, this implicitly assumes $X_{i}$ is convex, since otherwise $\rho(f_{i})=\infty$ and the bound is meaningless. When $X_{i}$ is nonconvex, an additional step is required to ‘primalize’ the solutions for $i\not\in\mathcal{J}$. We explore such primalization steps in Section 3.3. Moreover, the approach described in udell2016bounding requires optimizing over the solution set of the convex relaxation (CoP), which in some cases might be a difficult task. In contrast, the approach we present in Section 3 only assumes access to the value of $v^{*}$, which can be obtained by either solving the dual (D) or the convex relaxation (CoP), whichever one is easier to optimize.
2.4.2.   General domains
When dealing with nonconvex domains, we have $\operatorname{dom}f_{i}^{**}=\operatorname{conv}X_{i}\not=X_{i}$. In particular, $\rho(f_{i})=\infty$, and thus a more general notion of nonconvexity is needed. As in bertsekas2014constrained , let us define
$\gamma(f_{i}):=\sup_{x\in X_{i}}f_{i}(x)-\inf_{y\in X_{i}}f_{i}(y).$ (10)
In order to prove a duality gap bound similar to the one of Theorem 2.4.2, an additional assumption on the domain of the functions is necessary.
{restatable}
assumption AssNonConvexDomain For all $i=1,\dots,n$ and all $x_{i}\in\operatorname{conv}X_{i}$, there exists $\hat{x}_{i}\in X_{i}$ such that $A_{i}\hat{x}_{i}\leq A_{i}x_{i}$. We will discuss in Remark 1 the relevance of this assumption in real-world applications. For now, let us recall and prove the following theorem, which states that the duality gap can be bounded by a term depending on the number of constraints $m$ and on $\max_{i}\gamma(f_{i})$, but not on the number of functions $n$ aubin1976estimates ; bertsekas2014constrained .
{restatable}
thmNonConvexDomainDualityGap Let $v^{*}$ be the dual optimal value defined in (6) and $\gamma(f_{i})$ be defined in (10). Suppose that Assumption 2.4.2 holds. Then there exists $x^{*}\in\mathbb{R}^{d}$ such that $x^{*}_{i}\in X_{i}$ for all $i$, $Ax^{*}\leq b$ and
$v^{*}\leq\sum_{i=1}^{n}f_{i}(x_{i}^{*})\leq v^{*}+(m+1)\max_{i}\gamma(f_{i}).$
Proof
By Lemma 2.2, there exists $\bar{x}\in\mathbb{R}^{d}$ and set $S\subset\{1,\dots,n\}$ of cardinality at most $m+1$ such that $A\bar{x}\leq b$, $\bar{x}_{i}\in X_{i}$ for $i\not\in S$ and $\bar{x}_{i}\in\operatorname{conv}X_{i}$ such that
$v^{*}=\sum_{i\not\in S}f_{i}(\bar{x}_{i})+\sum_{i\in S}f_{i}^{**}(\bar{x}_{i}).$
In contrast with the proof of Theorem 2.4.1, $\bar{x}$ is not necessarily primal feasible since $X_{i}$ is not convex.
For $i\in S$, $\bar{x}_{i}\in\operatorname{conv}X_{i}$ and by Assumption 2.4.2 there exists $\hat{x}_{i}\in X_{i}$ such that $A_{i}\hat{x}_{i}\leq A_{i}\bar{x}_{i}$. Let then $x_{i}^{*}=\hat{x}_{i}\in X_{i}$.
For $i\not\in S$, let $x^{*}_{i}=\bar{x}_{i}$. We then have
$Ax^{*}=\sum_{i\in S}Ax^{*}_{i}+\sum_{i\not\in S}Ax^{*}_{i}\leq\sum_{i\in S}A\bar{x}_{i}+\sum_{i\not\in S}A\bar{x}_{i}\leq b,$
so that $x^{*}$ is primal feasible, and
$v^{*}$ $\leq\sum_{i=1}^{n}f_{i}(x_{i}^{*})$
$=\sum_{i\not\in S}f_{i}(x_{i}^{*})+\sum_{i\in S}f_{i}(x_{i}^{*})$
$=v^{*}+\sum_{i\in S}(f_{i}(x_{i}^{*})-f_{i}^{**}(\bar{x}_{i}))$
$\leq v^{*}+\sum_{i\in S}\sup_{x_{i}\in X_{i}}f_{i}(x_{i})-\inf_{x_{i}\in X_{i}}f_{i}(x_{i})$
$\leq v^{*}+(m+1)\max_{i}\gamma(f_{i}).$
The fact that $\inf_{x_{i}\in X_{i}}f_{i}(x_{i})\leq f_{i}^{**}(\bar{x}_{i})$ follows from the fact that the function $g(x_{i})=\inf_{y\in X_{i}}f_{i}(y)$ is a convex minorant of $f_{i}$ on $X_{i}$.
As for Theorem 2.4.1, this theorem is an existence result, and its proof does not yield a constructive algorithm. For MILPs whose bidual (CoP) has a unique solution, vujanic2014large designs an algorithm satisfying the bound of Theorem 2.4.2.
Remark 1
While Assumption 2.4.2 may seem quite restrictive, there exist several real-world applications for which it is satisfied. For example, in the unit commitment problem introduced in Section 1.1, having some $x_{i}\not\in X_{i}$ but $x_{i}\in\operatorname{conv}X_{i}$ means that plant $i$ should be producing between 0 and the minimum amount allowed if turned on. Obviously, this is not possible, but one can simply set the production to the minimum amount allowed, and this will only increase the total amount of electricity produced over all plants, and hence will not violate the demand constraint. The assumption also holds, for example, for supply chain problems vujanic2014large and portfolio optimization baumann2013portfolio .
If Assumption 2.4.2 does not hold, one can instead solve the following perturbed problem
$\begin{split}\min_{x_{i}\in\operatorname{conv}X_{i}}&\ \sum_{i=1}^{n}f_{i}^{**}(x_{i})\\ \text{subject to }&\ Ax\leq b-\theta.\end{split}$ ($\text{CoP}_{\theta}$)
The next theorem states that for a well chosen value of $\theta$, there exists a solution of ($\text{CoP}_{\theta}$) whose duality gap can be bounded by the sum of two terms, one depending on the number of constraints $m$, and the second one characterizing the feasibility of ($\text{CoP}_{\theta}$) and the perturbation parameter $\theta$. This result was proven for MILPs in vujanic2016decomposition , and here we extend it to general nonconvex problems. Recall that $(A_{i})_{k,:}$ corresponds to the $k$-th row of the matrix $A_{i}$.
{restatable}
thmNonConvexDomainDualityGapTwo Let $v^{*}$ be the dual optimal value as defined in (6) and $\gamma(f_{i})$ be as defined in (10). Let $\theta\in\mathbb{R}^{m}$ be defined as
$\theta_{k}=(m+1)\max_{i}\left(\max_{x_{i}\in X_{i}}((A_{i})_{k,:})^{\top}x_{i}-\min_{x_{i}\in X_{i}}((A_{i})_{k,:})^{\top}x_{i}\right),$
and suppose that there exists $\xi>0$ and $\hat{x}_{i}\in\operatorname{conv}X_{i}$ for all $i$ such that
$\sum_{i=1}^{n}A_{i}\hat{x}_{i}\leq b-\theta-n\xi\mathbf{1}_{m}.$
Then, there exists a solution $x^{*}\in\mathbb{R}^{d}$, such that $x^{*}_{i}\in X_{i}$, $Ax^{*}\leq b$, and
$v^{*}\leq\sum_{i=1}^{n}f_{i}(x_{i}^{*})\leq v^{*}+\left((m+1)+\frac{\left\lVert\theta\right\rVert_{\infty}}{\xi}\right)\max_{i}\gamma(f_{i}).$
Proof
Let $v^{*}_{\theta}$ be the optimal value of ($\text{CoP}_{\theta}$). Applying Lemma 2.2 to $v^{*}_{\theta}$, there exists a set $S\subset\{1,\dots,n\}$ of cardinality at most $m+1$ and $\bar{x}\in\mathbb{R}^{d}$ such that $A\bar{x}\leq b-\theta$, $\bar{x}_{i}\in X_{i}$ for $i\not\in S$ and $\bar{x}_{i}\in\operatorname{conv}(X_{i})$ for $i\in S$ such that
$v^{*}_{\theta}=\sum_{i\in S}f_{i}^{**}(\bar{x}_{i})+\sum_{i\not\in S}f_{i}(\bar{x}_{i})$
Define $x^{*}\in\mathbb{R}^{d}$ as $x^{*}_{i}=\bar{x}_{i}$ for $i\not\in S$ and $x^{*}_{i}$ be any point in $X_{i}$ for $i\in S$. We then have
$Ax^{*}$ $=\sum_{i\in S}A_{i}x^{*}_{i}+\sum_{i\not\in S}A_{i}x^{*}_{i}=\sum_{i\in S}A_{i}(x^{*}_{i}-\bar{x}_{i})+\sum_{i\in S}A_{i}\bar{x}_{i}+\sum_{i\not\in S}A_{i}\bar{x}_{i}$
$=\sum_{i\in S}A_{i}(x^{*}_{i}-\bar{x}_{i})+\sum_{i=1}^{n}A_{i}\bar{x}_{i}$
$\leq\theta+(b-\theta)=b$
Thus $x^{*}$ is primal feasible. As in Theorem 2.4.2, we can then show that
$\sum_{i=1}^{n}f_{i}(x_{i}^{*})\leq v^{*}_{\theta}+(m+1)\max_{i}\gamma(f_{i}).$
It remains to show that
$v^{*}_{\theta}\leq v^{*}+\frac{\left\lVert\theta\right\rVert_{\infty}}{\xi}\max_{i}\gamma(f_{i}).$
We mirror the proof of (vujanic2016decomposition, , Theorem 3.3), except that we make no assumption on the uniqueness of the solution, or on the linearity of the objective functions. Let $\Psi_{\theta}$ be defined as $\Psi$ in (5), except for the perturbed problem, and let $\lambda^{*}_{\theta}=\operatorname*{arg\,max}_{\lambda\geq 0}\Psi_{\theta}(\lambda)$. By (nedic2009approximate, , Lemma 1), $\hat{x}$ satisfies, for all $\lambda\geq 0$,
$\left\lVert\lambda^{*}_{\theta}\right\rVert_{1}$ $\leq\frac{1}{n\xi}\left(\sum_{i=1}^{n}f_{i}^{**}(\hat{x}_{i})-\Psi_{\theta}(\lambda)\right)$
$\leq\frac{1}{n\xi}\left(\sum_{i=1}^{n}f_{i}^{**}(\hat{x}_{i})-\sum_{i=1}^{n}\inf_{x_{i}\in X_{i}}(f_{i}(x)+\lambda^{\top}A_{i}x_{i})-\lambda^{\top}(b-\theta)\right).$
Plugging $\lambda=0$, and because $f_{i}^{**}\leq f_{i}$, this gives
$\left\lVert\lambda^{*}_{\theta}\right\rVert_{1}$ $\leq\frac{1}{\xi}\max_{i}\gamma(f_{i}).$
Now consider ($\text{CoP}_{\theta}$) as an unperturbed program, and (CoP) as its perturbed version (perturbed by $+\theta$). Sensitivity analysis as in (boyd2004convex, , Section 5.6.2) tells us that since strong duality holds,
$v^{*}_{\theta}\leq v^{*}+(\lambda^{*}_{\theta})^{\top}\theta\leq v^{*}+\left\lVert\lambda^{*}_{\theta}\right\rVert_{1}\left\lVert\theta\right\rVert_{\infty}\leq v^{*}+\frac{\left\lVert\theta\right\rVert_{\infty}}{\xi}\max_{i}\gamma(f_{i}).$
As already discussed, the proofs of the duality gap bounds presented in Theorem 2.4.1, Theorem 2.4.2 and Theorem 1 are not constructive. The next section develops an approach to build solutions obtaining those bounds up to an arbitrarily small term.
3.   Solving nonconvex separable problems
We now come back to our original task of finding approximate solutions to problem (P). Recall that $v^{*}$ is the optimal value of the dual (D).
3.1.   Obtaining primal points
Let us define $z^{*}\in\mathbb{R}^{1+m}$ as
$z^{*}=\begin{pmatrix}v^{*}\\ b\end{pmatrix}.$ (11)
Since $v^{*}$ is the optimal value of (CoP), there exists $x^{*}\in\mathbb{R}^{d}$ such that $Ax^{*}\leq b$ and $v^{*}=\sum_{i=1}^{n}f_{i}^{**}(x_{i}^{*})$. In other words,
$z^{*}\in\sum_{i=1}^{n}\left\{\begin{pmatrix}f_{i}^{**}(x_{i})\\ A_{i}x_{i}\end{pmatrix}\mid x_{i}\in\operatorname{conv}X_{i}\right\}+\mathbb{R}^{m+1}_{+}.$ (12)
Let us define the corresponding sets
$\mathcal{C}_{i}$ =$=\left\{\begin{pmatrix}f_{i}^{**}(x_{i})\\ A_{i}x_{i}\end{pmatrix}\mid x_{i}\in\operatorname{conv}X_{i}\right\},\quad i=1,\dots,n,$
$\mathcal{C}$ $=\sum_{i=1}^{n}\mathcal{C}_{i}.$
We aim to approximate $z^{*}$ with elements of $\mathcal{C}_{i}$ where $f_{i}$ and $f_{i}^{**}$ match. The next lemma, proved in Appendix A, states that this is the case at extreme points of $\mathcal{C}_{i}$. {restatable}lemExtremePointsMatching Let $\tilde{x}_{i}\in\operatorname{conv}X_{i}$ such that $(f_{i}^{**}(\tilde{x}_{i}),A_{i}\tilde{x}_{i})$ is an extreme point of $\mathcal{C}_{i}$. Then $\tilde{x}_{i}\in X_{i}$ and $f_{i}(\tilde{x}_{i})=f_{i}^{**}(\tilde{x}_{i})$. To extract extreme points of $\mathcal{C}_{i}$, conditional gradient methods are strong candidates. At first sight, a natural objective function is the quadratic loss, giving rise to the following optimization problem
$\begin{split}\min_{z}\ &\frac{1}{2}\left\lVert z-z^{*}\right\rVert^{2}\ \text{ subject to}\ z\in\mathcal{C}+\mathbb{R}^{m+1}_{+}.\end{split}$ (13)
However, linear minimization on the unbounded set $\mathcal{C}+\mathbb{R}^{m+1}_{+}$, which is required for conditional gradient methods, is not achievable. Alternatively, one can solve
$\begin{split}\min_{z}\ &\frac{1}{2}\left\lVert z-z^{*}\right\rVert_{+}^{2}\ \text{ subject to}\ z\in\mathcal{C}.\end{split}$ (14)
Lemma 3.1
Problems (13) and (14) are equivalent.
Proof
It is quick to check that for any $c\in\mathcal{C}$,
$\left\lVert c-z^{*}\right\rVert_{+}^{2}=\inf_{\delta\in\mathbb{R}^{m+1}_{+}}\left\lVert c+\delta-z^{*}\right\rVert^{2}.$
Hence
$\inf_{c\in\mathcal{C}}\left\{\left\lVert c-z^{*}\right\rVert^{2}_{+}\right\}=\inf_{c\in\mathcal{C}}\inf_{\delta\in\mathbb{R}^{m+1}_{+}}\left\{\left\lVert c+\delta-z^{*}\right\rVert^{2}\right\}=\inf_{z\in\mathcal{C}+\mathbb{R}^{m+1}_{+}}\left\{\left\lVert z-z^{*}\right\rVert^{2}\right\}.$
The next lemma shows that any solution of (14) has optimal value 00 and yields a solution of (CoP). {restatable}lemEquivalenceBetweenFWObjectives Suppose $\tilde{z}=\begin{pmatrix}\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})\\ A\tilde{x}\end{pmatrix}$ is a solution of (14).
Then $\left\lVert\tilde{z}-z^{*}\right\rVert_{+}=0$ and $\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})=v^{*}$ and $A\tilde{x}\leq b$.
Proof
Let $x^{*}$ be a solution to the bi-dual problem (CoP). We then have
$z^{*}=\begin{pmatrix}v^{*}\\ b\end{pmatrix}=\begin{pmatrix}\sum_{i=1}^{n}f_{i}^{**}(x_{i}^{*})\\ Ax^{*}\end{pmatrix}+\begin{pmatrix}0\\ b-Ax^{*}\end{pmatrix}\in\sum_{i=1}^{n}\mathcal{C}_{i}+\mathbb{R}^{m+1}_{+}.$
In particular, we have
$\frac{1}{2}\left\lVert\begin{pmatrix}\sum_{i=1}^{n}f_{i}^{**}(x_{i}^{*})\\ Ax^{*}\end{pmatrix}-z^{*}\right\rVert_{+}^{2}=\frac{1}{2}\left\lVert\begin{pmatrix}0\\ Ax^{*}-b\end{pmatrix}\right\rVert_{+}^{2}=0,$
since $Ax^{*}-b\leq 0$. This implies that the minimum value of (14) is 0. In particular, if $\begin{pmatrix}\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})\\ A\tilde{x}\end{pmatrix}$ is a solution of (14), then $A\tilde{x}\leq b$ and $\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})\leq v^{*}$. Since $v^{*}$ is the optimal value of (CoP), we actually have $\sum_{i=1}^{n}f_{i}^{**}(\tilde{x}_{i})=v^{*}$.
We use the classical Frank Wolfe algorithm to solve (14). The algorithm starts from some guess $z_{0}\in\mathcal{C}$. At each iteration $k$, the gradient of the objective function with respect to $z^{k}$ is given by $(\alpha_{k},g^{k})\in\mathbb{R}\times\mathbb{R}^{m}$ with $\alpha_{k}=\max(z^{k}_{1}-v^{*},0)$ and $g^{k}=\max(z^{k}_{2:m+1}-b,0)$. The linear minimization step then reads
$s^{k}\in\operatorname*{arg\,min}_{z\in\sum_{i=1}^{n}\mathcal{C}_{i}}z^{\top}(\alpha_{k},g^{k}).$ (15)
This amounts to solving, for each $i\in\{1,\dots,n\}$,
$y_{i}^{k}\in\operatorname*{arg\,min}_{x_{i}\in\operatorname{conv}X_{i}}\alpha_{k}f_{i}^{**}(x_{i})+(g^{k})^{\top}A_{i}x_{i},$ (16)
and setting $s^{k}=\sum_{i=1}^{n}y_{i}^{k}$. If $\alpha_{k}=0$, the above is a linear minimization over $\operatorname{conv}X_{i}$. Otherwise, $\alpha_{k}>0$, and we can can rewrite (16) as
$y_{i}^{k}\in\operatorname*{arg\,max}_{x_{i}\in\operatorname{conv}X_{i}}\left(-\frac{A_{i}^{\top}g^{k}}{\alpha_{k}}\right)^{\top}x_{i}-f_{i}^{**}(x_{i})=\partial f_{i}^{*}\left(-\frac{A_{i}^{\top}g^{k}}{\alpha_{k}}\right).$
The last few lines are central to our approach. They tell us that the key subproblem in conditional gradient methods, namely the linear minimization oracle (15), is as tractable as computing the conjugate of the functions $f_{i}$.
Remark 2
It is interesting to look into the computation of (16) when the function $f_{i}$ is concave and its domain $X_{i}$ is a polytope. In that case, the convex envelope of $f_{i}$ on $X_{i}$ is vertex polyhedral, meaning that it is equal to the convex envelope of $f_{i}$ restricted to the vertices of $X_{i}$ falk1976successive . Importantly here, the minimum of the convex envelope of such functions, and therefore also the computation of (16), then comes down to solving a linear system and can thus be obtained efficiently tardella2004existence . Note that functions with a vertex polyhedral convex envelope are not restricted to concave functions but also include, for example, the more general edge-concave functions tardella2004existence and multilinear functions on certain polytopes rikun1997convex .
We summarize our full approach in Algorithm 2. Note that while there always exists an extreme point of $\mathcal{C}_{i}$ which solves (16), a non-extreme point might yield the same objective value. Without loss of generality, we assume that we always get an extreme point, and hence $y_{i}^{k}\in X_{i}$ and $f_{i}(y_{i}^{k})=f_{i}^{**}(y_{i}^{k})$. In the applications we consider in section Section 5, this is never an issue.
At any iteration $k$ of the algorithm, the suboptimality gap $\left\lVert z^{k}-z^{*}\right\rVert_{+}$ can be used as stopping criterion. Alternatively, the quality of the approximation can be bounded using classical convergence results on the Frank-Wolfe method.
Theorem 3.1
Suppose we run Algorithm 2 for $K$ iterations with $\eta_{k}=\frac{2}{k+2}$ or $\eta_{k}=\min\left(1,\frac{(v^{k})^{\top}(z^{k}-s^{k})}{\left\lVert z^{k}-s^{k}\right\rVert^{2}}\right)$, where $v^{k}=\max(z^{k}-z^{*},0)$. Then the output satisfies
$\left\lVert z^{K}-z^{*}\right\rVert^{2}_{+}\leq\frac{4D_{\mathcal{C}}^{2}}{K+1},$
where $D_{\mathcal{C}}=\max_{z,z^{\prime}\in\mathcal{C}}\left\lVert z-z^{\prime}\right\rVert$.
Proof
This is a direct consequence of the analysis of Frank-Wolfe algorithm, see for example (jaggi2011sparse, , Theorem 2.3) for the stepsize $\eta_{k}=\frac{2}{k+2}$ or pedregosa2020linearly for the other choice of $\eta_{k}$. Note that the feasible set $\mathcal{C}$ we are considering is not necessarily convex, and hence it might be that $z^{K}\not\in\mathcal{C}$. This is however not an issue for the presented bound to hold.
After running Algorithm 2 for $K$ iterations, the output $z^{K}$ is a convex combination of the vectors $\{s_{k}\}_{k=0}^{K-1}$. In other words, for $\gamma\in\Delta_{K}$, we have
$z^{K}=\sum_{k=0}^{K-1}\gamma_{k}s_{k}=\sum_{k=0}^{K-1}\gamma_{k}\sum_{i=1}^{n}\begin{pmatrix}f_{i}(y_{i}^{k})\\ A_{i}y_{i}^{k}\end{pmatrix}.$ (17)
Algorithm 2 Frank Wolfe algorithm to obtain primal points
1:   Input: $z^{*}\in\mathbb{R}^{1+m}$, number of iterations $K$.
2:  For all $i=1,\dots,n$, choose $y_{i}^{0}\in X_{i}$ such that $f_{i}(y_{i}^{0})=f_{i}^{**}(y_{i}^{0})$.
3:  Set $z^{0}=\sum_{i=1}^{n}\begin{pmatrix}f_{i}(y_{i}^{0})\\ A_{i}y_{i}^{0}\end{pmatrix}$
4:  for $k=0,\dots,K-1$ do
5:     Set $\alpha_{k}=\max(z^{k}_{1}-v^{*},0)$ and $g^{k}=\max(z^{k}_{2:m+1}-b,0)$.
6:     for $i=1,\dots,n$ do
7:        if $\alpha_{k}=0$ then
$y_{i}^{k}\in\operatorname*{arg\,min}_{x_{i}\in\operatorname{conv}X_{i}}(A_{i}^{\top}g^{k})^{\top}x_{i}$
8:        else
$y_{i}^{k}\in\partial f_{i}^{*}\left(-\frac{A_{i}^{\top}g^{k}}{\alpha_{k}}\right)$
9:     end for
10:     Set $s^{k}=\sum_{i=1}^{n}\begin{pmatrix}f_{i}(y_{i}^{k})\\ A_{i}y_{i}^{k}\end{pmatrix}$.
11:     Pick some stepsize $\eta_{k}>0$.
12:     $z^{k+1}=(1-\eta_{k})z_{k}+\eta_{k}s_{k}$.
13:  end for
14:  return  $z^{K}$
Remark 3
Algorithm 2 optimizes a (non-strongly) convex, smooth objective function. In this scenario, Frank-Wolfe algorithms cannot achieve better than $O\left(\frac{1}{K}\right)$ convergence rate lan2013complexity . Therefore, another conditional gradient method would not achieve better performance than Algorithm 2, at least in theory. Of course, one is free to implement any variant of Frank-Wolfe, for example pairwise mitchell1974finding ; lacoste2015global , away-steps wolfe1970convergence , or classical Frank-Wolfe with a backtracking linesearch pedregosa2020linearly . The only requirement is that it must be possible to keep track of the coefficients in the convex representation of the final result $z^{K}$.
3.2.   Trimming the solution
Let us define $w^{K}\in\mathbb{R}^{1+m+n}$ and $\mathcal{A}_{K}\subset\mathbb{R}^{1+m+n}$ as
$w^{K}=\begin{pmatrix}z^{K}\\ \mathbf{1}_{n}\end{pmatrix},\ \text{ and }\ \mathcal{A}_{K}=\left\{\begin{pmatrix}f_{i}(y_{i}^{k})\\ A_{i}y_{i}^{k}\\ e_{i}\end{pmatrix}\mid k=0,\dots,K-1,i=1,\dots n\right\}.$ (18)
From Equation 17, we get that
$w^{K}=\sum_{k=0}^{K-1}\gamma_{k}\sum_{i=1}^{n}\begin{pmatrix}f_{i}(y_{i}^{k})\\ A_{i}y_{i}^{k}\\ e_{i}\end{pmatrix},$ (19)
where $\{e_{i}\}_{i=1}^{n}$ is the canonical basis of $\mathbb{R}^{n}$. In other words, $w^{K}\in\operatorname{Po}(\mathcal{A}_{K})$.
We can thus apply Algorithm 1 with $w^{*}=w^{K}$. From Theorem 2.2, this yields a vector $\alpha\in\mathbb{R}^{1+m+n}_{+}$ such that
$w^{K}=\sum_{l=1}^{n+m+1}\alpha_{l}\begin{pmatrix}f_{i_{l}}(y_{i_{l}}^{k_{l}})\\ A_{i_{l}}y_{i_{l}}^{k_{l}}\\ e_{i_{l}}\end{pmatrix}=\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}\begin{pmatrix}f_{i}(y_{i}^{k_{l}})\\ A_{i}y_{i}^{k_{l}}\\ e_{i}\end{pmatrix},$ (20)
where we regrouped the terms in $i$ by defining $I_{i}:=\{l\mid i_{l}=i\}$ for each $i=1,\dots,n$. This yields the following theorem.
Proposition 3.1
Suppose we run Algorithm 1 with input $\mathcal{A}_{K}$ and $w^{K}$ defined in (19). It outputs a vector $\alpha\in\mathbb{R}^{n+m+1}_{+}$ such that
$w^{K}=\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}\begin{pmatrix}f_{i}(y_{i}^{k_{l}})\\ A_{i}y_{i}^{k_{l}}\\ e_{i}\end{pmatrix}.$ (21)
In particular, for all $i$, we have $\sum_{l\in I_{i}}\alpha_{l}=1$. Moreover, $|I_{i}|>1$ for at most $m+1$ indices $i$.
Proof
The fact that $\sum_{l\in I_{i}}\alpha_{l}=1$ follows from the last $n$ coordinates of $w^{K}$ being equal to $1$. In particular, we must have $|I_{i}|\geq 1$ for all $i$. Since $\sum_{i=1}^{n}|I_{i}|=n+m+1$, there can only be at most $m+1$ indices $i$ such that $|I_{i}|>1$.
Remark 4
If Algorithm 1 is applied with a dense QR decomposition and column addition/deletion as described before Lemma 2.1, the complexity is $O(nK(n+m)^{2})$. Alternatively, since the vectors in $\mathcal{A}_{K}$ are sparse (only one coordinate among the last $n$ is nonzero), sparse solvers might be more efficient in practice.
3.3.   Final reconstruction
In this section, we explore different ways to reconstruct a solution of the primal problem (P) from the decomposition (21). We start with the following lemma, which is straightforward from the convergence rate derived in Theorem 3.1. The reason we split the coordinates is that it will be useful to differentiate between the coordinates related to function values and the coordinates related to the affine constraints later on.
Lemma 3.2
Let $v^{*}$ be the optimal value of (D) and (CoP). Assume that we run Algorithm 2 for $K$ iterations, followed by Algorithm 1 to trim the number of elements. Then
$w^{K}_{1}\leq v^{*}+\frac{2D_{\mathcal{C}}}{\sqrt{K+1}},$
$\left\lVert w^{K}_{2:m+1}-b\right\rVert_{+}\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}.$
Proof
Since $z^{K}$, namely the output of Algorithm 2, consists of the first $1+m$ coordinates of $w^{K}$, we have in particular $w^{K}_{1}=z^{K}_{1}$ and $w^{K}_{2:m+1}=z^{K}_{2:m+1}$. The result then follows from the convergence rate from Theorem 3.1.
3.3.1.   Convex domains
When the function domains are convex, a natural candidate solution is $\bar{x}\in\mathbb{R}^{d}$ defined as
$\bar{x}_{i}=\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}},\quad i=1,\dots,n.$ (22)
The next result, which we consider to be the main result of this work, recovers the bound of Theorem 2.4.1 for the proposed solution (22), up to an additional term which decreases with $K$.
Theorem 3.2 (Main result)
Suppose that $X_{i}$ is convex for all $i=1,\dots,n$, and let $v^{*}$ be the optimal value of (D) and (CoP). Assume that we run Algorithm 2 for $K$ iterations, followed by Algorithm 1 to trim the number of elements. Let $\bar{x}\in\mathbb{R}^{d}$ be the final solution defined in (22). Then $\bar{x}$ satisfies
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})$ $\leq v^{*}+\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}+(m+1)\max_{i}\rho(f_{i}),$
$\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-b\right\rVert_{+}$ $\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}.$
Proof
Observe that
$\begin{split}\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-v^{*}&=\left(\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}\right)+\left(w^{K}_{1}-v^{*}\right)\\ \left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-b\right\rVert_{+}&\leq\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}\right\rVert_{+}+\left\lVert w^{K}_{2:m+1}-b\right\rVert_{+}\end{split}$ (23)
We can apply Lemma 3.2 to bound the last term of each equation. Moreover,
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}$ $=\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})$
$=\sum_{i,|I_{i}|>1}\left(f_{i}\left(\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}}\right)-\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})\right)$
$\leq\sum_{i,|I_{i}|>1}\rho(f_{i})$
$\leq(m+1)\max_{i}\rho(f_{i}),$
and
$\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}$ $=\sum_{i=1}^{n}A_{i}\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}}-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}A_{i}y_{i}^{k_{l}}=0.$
3.3.2.   General domains
When the domains $X_{i}$ are not necessarily convex, the solution $\bar{x}$ defined in (22) is in general not feasible, as for $i$ such that $|I_{i}|>1$, $\bar{x}_{i}$ is a non-trivial convex combination of elements of $X_{i}$. In this section, we suggest several approaches to construct feasible solutions.
3.3.2.1.   Reconstruction via sampling
Our first reconstruction scheme is a randomized approach. We suggest setting the solution $\bar{x}_{i}$ as $y_{i}^{k_{l}}$ with probability $\alpha_{l}$. In other words, for each $i=1,\dots,n$, define $L_{i}$ as the random variable taking values in $I_{i}$ such that for all $l\in I_{i}$, $P(L_{i}=l)=\alpha_{l}$. The reconstruction scheme is then
$\begin{split}&l_{i}\sim L_{i}\\ &\bar{x}_{i}=y_{i}^{k_{l_{i}}}\end{split}\quad i=1,\dots,n.$ (24)
Since $y_{i}^{k_{l}}\in X_{i}$ for all $i$ and $l\in I_{i}$, we immediately get that $\bar{x}_{i}\in X_{i}$ for all $i$. We also get the following duality gap bound on the obtained solutions. In particular, the obtained solution satisfies the constraint in expectation (up to an arbitrarily small term).
Theorem 3.3
Let $v^{*}$ be the optimal value of (D) and (CoP). Assume that we run Algorithm 2 for $K$ iterations, followed by Algorithm 1 to trim the number of elements. Let $\bar{x}\in\mathbb{R}^{d}$ be the final solution defined in (24). Then $\bar{x}$ satisfies
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})$ $\leq v^{*}+\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}+(m+1)\max_{i}\gamma(f_{i}),$
$\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\bar{x}_{i}-b\right]$ $\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}.$
Proof
Observe that
$\begin{split}\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-v^{*}&=\left(\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}\right)+\left(w^{K}_{1}-v^{*}\right)\\ \sum_{i=1}^{n}A_{i}\bar{x}_{i}-b&=\left(\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}\right)+\left(w^{K}_{2:m+1}-b\right)\\ &\leq\left(\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}\right)+\left\lVert w^{K}_{2:m+1}-b\right\rVert_{+}.\end{split}$ (25)
Lemma 3.2 gives a bound on the last term of each equation. Moreover, we have
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}$ $=\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})$
$=\sum_{i,|I_{i}|>1}\left(f_{i}(\bar{x}_{i})-\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})\right)$
$\leq\sum_{i,|I_{i}|>1}\gamma(f_{i})$
$\leq(m+1)\max_{i}\gamma(f_{i}),$
and
$\mathbb{E}\left[\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}\right]$ $=\sum_{i=1}^{n}A_{i}\mathbb{E}\left[\bar{x}_{i}\right]-w^{K}_{2:m+1}$
$=\sum_{i=1}^{n}A_{i}\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}}-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}A_{i}y_{i}^{k_{l}}=0.$
3.3.2.2.   Reconstruction under additional domain assumption
The above reconstruction gives a solution which is feasible in expectation. In particular, one can build solutions of the form (24) by repeatedly sampling until a good enough solution is obtained. However, there is no guarantee that a fully feasible solution will be obtained with that procedure. In order to construct such a solution, we require additional assumptions on the feasible set.
We already explored one common such assumption in Section 2.4, namely Assumption 2.4.2. We restate it here for completeness.
[…]
* Recall that we discussed in Remark 1 several real-world applications for which this assumption holds. We can then construct a final solution as follows. If index $i$ is such that $|I_{i}|>1$, then by Assumption 2.4.2 there exists $\hat{x}_{i}\in X_{i}$ such that
$A_{i}\hat{x}_{i}\leq A_{i}\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}}$ (26)
We can then build a final solution as
$\begin{split}\bar{x}_{i}=\begin{cases}\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}}&\text{ if }|I_{i}|=1,\\ \hat{x}_{i}&\text{ if }|I_{i}|>1.\end{cases}\end{split}$ (27)
The next theorem states that the resulting solution recovers the duality gap bound of Theorem 2.4.2, up to an additional term decreasing with $K$.
{restatable}
thmNonConvexDomainConvergenceOne Let $v^{*}$ be the optimal value of (D) and (CoP) and suppose that Assumption 2.4.2 holds. Assume that we run Algorithm 2 for $K$ iterations, followed by Algorithm 1 to trim the number of elements. Let $\bar{x}\in\mathbb{R}^{d}$ be the final solution defined in (27). Then $\bar{x}$ satisfies
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})$ $\leq v^{*}+\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}+(m+1)\max_{i}\gamma(f_{i}),$
$\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-b\right\rVert_{+}$ $\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}.$
Proof
Observe that
$\begin{split}\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-v^{*}&=\left(\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}\right)+\left(w^{K}_{1}-v^{*}\right),\\ \left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-b\right\rVert_{+}&\leq\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}\right\rVert_{+}+\left\lVert w^{K}_{2:m+1}-b\right\rVert_{+}.\end{split}$ (28)
We can apply Lemma 3.2 to bound the last term of each equation. Moreover, as in the proof of Theorem 3.3, we have
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}\leq(m+1)\max_{i}\gamma(f_{i}).$
Now,
$\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}\right\rVert_{+}$ $=\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}A_{i}y_{i}^{k_{l}}\right\rVert_{+}$
$=\left\lVert\sum_{i,|I_{i}|>1}\left(A_{i}\hat{x}_{i}-\sum_{l\in I_{i}}\alpha_{l}A_{i}y_{i}^{k_{l}}\right)\right\rVert_{+}$
$=0,$
where the last equality is a consequence of (26).
3.3.2.3.   Reconstruction after solving perturbed problem
Assumption 2.4.2 can sometimes be restrictive. In such cases, following vujanic2016decomposition , we consider a perturbed version of (CoP), namely ($\text{CoP}_{\theta}$), with $\theta\in\mathbb{R}^{m}$ defined as
$\theta_{k}:=(m+1)\max_{i}\left(\max_{x_{i}\in X_{i}}((A_{i})_{k,:})^{\top}x_{i}-\min_{x_{i}\in X_{i}}((A_{i})_{k,:})^{\top}x_{i}\right).$ (29)
This will ensure that the error made when constructing a final solution is small enough so that the solution is still primal feasible. We also need the following technical assumption, which is a Slater-type condition on the perturbed problem ($\text{CoP}_{\theta}$).
Assumption 3.1
For all $i=1,\dots,n$, there exists $\xi>0$ and $\hat{x}_{i}\in\operatorname{conv}X_{i}$ such that
$\sum_{i=1}^{n}A_{i}\hat{x}_{i}\leq b-\theta-n\xi\mathbf{1}_{m}.$
Finally, we suggest the following reconstruction scheme
$\begin{split}l_{i}&=\operatorname*{arg\,max}_{l\in I_{i}}\alpha_{l},\\ \bar{x}_{i}&=y_{i}^{k_{l_{i}}},\end{split}\quad i=1,\dots,n.$ (30)
Taking the maximum over all coefficients in the convex combinations allows us to prove a stronger bound than in Theorem 1.
{restatable}
thmNonConvexDomainConvergenceTwo Let $v^{*}$ be the optimal value of (D) and (CoP) and suppose that Assumption 3.1 holds. Suppose that we apply the method of Sections 3.1 and 3.2 to problem ($\text{CoP}_{\theta}$) with $\theta$ defined in (29). Assume that we run Algorithm 2 for $K$ iterations, followed by Algorithm 1 to trim the number of elements. Let $\bar{x}\in\mathbb{R}^{d}$ be the final solution defined in (30). Finally, let $q=|\{i\in\{1,\dots,n\}\mid|I_{i}|>1\}|$ be the number of indices for which the output of the Carathédory algorithm gives a non-trivial convex combination. Then $\bar{x}$ satisfies
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})$ $\leq v^{*}+\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}+\left(\frac{q(m+1)}{q+(m+1)}+\frac{\left\lVert\theta\right\rVert_{\infty}}{\xi}\right)\max_{i}\gamma(f_{i}),$
$\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-b\right\rVert_{+}$ $\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}.$
Proof
Let $v^{*}_{\theta}$ be the optimal value of ($\text{CoP}_{\theta}$). Observe that
$\begin{split}\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-v^{*}&=\left(\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}\right)+\left(w^{K}_{1}-v^{*}_{\theta}\right)+\left(v^{*}_{\theta}-v^{*}\right)\\ \sum_{i=1}^{n}A_{i}\bar{x}_{i}-b&\leq\left(\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}\right)-\theta+\left(w^{K}_{2:m+1}-(b-\theta)\right)\\ &\leq\left(\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}\right)-\theta+\mathbf{1}_{m}\left\lVert w^{K}_{2:m+1}-(b-\theta)\right\rVert_{+}\end{split}$ (31)
Since we assumed that we applied the method to the perturbed problem ($\text{CoP}_{\theta}$), we can apply Lemma 3.2 to get
$\left(w^{K}_{1}-v^{*}_{\theta}\right)$ $\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}},$
$\left\lVert w^{K}_{2:m+1}-(b-\theta)\right\rVert_{+}$ $\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}.$
Now,
$\sum_{i=1}^{n}A_{i}\bar{x}_{i}-w^{K}_{2:m+1}$ $=\sum_{i=1}^{n}A_{i}\bar{x}_{i}-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}A_{i}y_{i}^{k_{l}}$
$=\sum_{i,|I_{i}|>1}\left(A_{i}\bar{x}_{i}-\sum_{l\in I_{i}}\alpha_{l}A_{i}y_{i}^{k_{l}}\right)$
$\leq\theta.$
This proves the inequalities on the feasibility. Moreover, we showed in Theorem 1 that
$v^{*}_{\theta}\leq v^{*}+\frac{\left\lVert\theta\right\rVert_{\infty}}{\xi}\max_{i}\gamma(f_{i}).$
Therefore, it only remains to show that
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}\leq\frac{q(m+1)}{q+(m+1)}\max_{i}\gamma(f_{i}).$
Indeed, we have
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-w^{K}_{1}$ $=\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})$
$=\sum_{i,|I_{i}|>1}\left(f_{i}(y_{i}^{k_{l_{i}}})-\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})\right)$
=$\sum_{i,|I_{i}|>1}\left((1-\alpha_{l_{i}})f_{i}(y_{i}^{k_{l_{i}}})-\sum_{l\in I_{i},l\not=l_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})\right)$
=$\sum_{i,|I_{i}|>1}(1-\alpha_{l_{i}})\left(f_{i}(y_{i}^{k_{l_{i}}})-\sum_{l\in I_{i},l\not=l_{i}}\frac{\alpha_{l}}{1-\alpha_{l_{i}}}f_{i}(y_{i}^{k_{l}})\right)$
$\leq\max_{i}\gamma(f_{i})\sum_{i,|I_{i}|>1}(1-\alpha_{l_{i}}).$
We now show that
$\sum_{i,|I_{i}|>1}(1-\alpha_{l_{i}})\leq\frac{q(m+1)}{q+(m+1)}.$
First observe that since $\alpha_{l_{i}}=\max_{l\in I_{i}}\alpha_{l}$ and $\sum_{l\in I_{i}}\alpha_{l}=1$, we must have $\alpha_{l_{i}}\geq\frac{1}{|I_{i}|}$. Moreover we have $\sum_{i,|I_{i}|>1}(|I_{i}|-1)\leq(m+1)$. We therefore look for an upper bound on the program:
$\max_{u_{i}\geq 2}\quad$ $\sum_{i=1}^{q}(1-\frac{1}{u_{i}})$
s.t. $\sum_{i=1}^{q}(u_{i}-1)\leq m+1.$
Since the program is concave and invariant under permutation, the solution must have all $u_{i}$’s equal so that it can be rewritten
$\max_{u\geq 2}\quad$ $q(1-\frac{1}{u})$
s.t. $q(u-1)\leq m+1.$
The Lagrangian reads
$L(u,\lambda,\mu)=q(1-\frac{1}{u})+\lambda(u-2)+\mu\left(m+1-q(u-1)\right).$
Maximization with respect to $u$ gives
$\frac{q}{u^{2}}+\lambda-\mu q=0\Rightarrow\frac{1}{u}=\sqrt{\mu-\lambda/q},$
and the dual then reads
$\min_{\mu\geq\lambda/q\geq 0}\quad$ $q-q\sqrt{\mu-\lambda/q}-2\lambda+\mu(m+1)+\mu q-q\sqrt{\mu-\lambda/q}.$
By the change of variable $\gamma=\sqrt{\mu-\lambda/q}$, we get
$\min_{\gamma,\lambda\geq 0}\quad$ $q-2q\gamma-2\lambda+(\gamma^{2}+\frac{\lambda}{q})(m+1)+(\gamma^{2}+\frac{\lambda}{q})q$
First order optimality gives $\lambda=0$ (since $q\leq m+1$) and $\gamma=\frac{q}{q+(m+1)}$. The corresponding optimal value is then $\frac{q(m+1)}{q+(m+1)}$.
In this section, we saw that $q\leq m+1$. Plugging in $q=(m+1)$ in the above theorem gives a term $\frac{m+1}{2}$ in the bound, which improves the term in Theorem 1 by a factor 2.
Remark 5
The approach presented in this section constructs solutions which meet the duality gap bounds presented in Section 2.4, up to an additional term $O(1/\sqrt{K})$, both in terms of functions values and feasibility. We argue that this not a major limitation. Indeed, in order to obtain a fully feasible solution, in practice one can solve a slightly perturbed version of the original problem. We describe one such way in the experimental section. Moreover, the additional term $O(1/\sqrt{K})$ in the function value bounds is typically negligible compared to the term in $(m+1)\max_{i}\rho(f_{i})$ or $(m+1)\max_{i}\gamma(f_{i})$.
4.   Extension to approximate Carathéodory representations
In Section 3.2, we showed how one can trim the result of the output of the first Frank-Wolfe algorithm by applying Algorithm 1, which computes an exact conic Carathéodory representation of $w^{K}$. In this section, we show that one can also compute cheaper approximate Carathéodory representations of $w^{K}$. Those approximate representations can be obtained by running a conditional gradient method. Those methods have the advantage of reducing the dependency on $m$ in the duality gap bound, at the expense of an exponentially decreasing error term which depends on the number of iterations of the conditional gradient method.
First let us recall that, after running Algorithm 2, we obtain $z^{K}\in\mathbb{R}^{1+m}$ with
$z^{K}=\sum_{k=0}^{K-1}\gamma_{k}\sum_{i=1}^{n}\begin{pmatrix}f_{i}(y_{i}^{k})\\ A_{i}y_{i}^{k}\end{pmatrix},$
where $\sum_{k=0}^{K-1}\gamma_{k}=1$. Moreover, recall from (18) that $w^{K}$ and $\mathcal{A}_{K}$ are defined as
$w^{K}=\begin{pmatrix}z^{K}\\ \mathbf{1}_{n}\end{pmatrix}=\sum_{k=0}^{K-1}\gamma_{k}\sum_{i=1}^{n}\begin{pmatrix}f_{i}(y_{i}^{k})\\ A_{i}y_{i}^{k}\\ e_{i}\end{pmatrix},$ (32)
$\mathcal{A}_{K}=\left\{\begin{pmatrix}f_{i}(y_{i}^{k})\\ A_{i}y_{i}^{k}\\ e_{i}\end{pmatrix}\mid k=0,\dots,K-1,i=1,\dots n\right\},$ (33)
In particular, we have $w^{K}/n\in\operatorname{conv}\mathcal{A}_{K}$. Instead of computing an exact Carathéodory decomposition of $w^{K}/n$, we look for an approximate decomposition, namely a set of vectors whose convex combination is a good approximation of $w^{K}$. Following on ideas from combettes2023revisiting , we therefore suggest applying a Frank-Wolfe algorithm to the following problem
$\operatorname*{arg\,min}_{w\in\operatorname{conv}\mathcal{A}_{K}}h(w):=\frac{1}{2}\left\lVert w-\frac{w^{K}}{n}\right\rVert^{2}.$ (34)
The reason behind using the conditional gradient method is that the output is a convex combination of vectors in $\mathcal{A}_{K}$, and therefore an approximate Carathéodory representation of $w^{K}/n$. The approximation error will be related to the convergence rate of the conditional gradient method. In this section, we will focus on the fully-corrective Frank-Wolfe (FCFW) algorithm, as it gives the best theoretical results combettes2023revisiting , but we will see how one can extend to other conditional gradient methods at the end of the section. The FCFW method starts with some $w^{K,0}\in\mathcal{A}_{K}$, sets $s^{0}=w^{K,0}$ and defines an active set $\mathcal{S}^{0}=\{s^{0}\}$. At each iteration $t\geq 1$, given some $w^{K,t-1}$ and an active set $\mathcal{S}^{t-1}$, the method makes a call to the linear minimization oracle
$s^{t}\in\operatorname*{arg\,min}_{s\in\operatorname{conv}(\mathcal{A}_{K})}s^{\top}\nabla h(w^{K,t-1}).$ (35)
Since this is a linear objective, the minimizer is attained at an extreme point of $\operatorname{conv}(\mathcal{A}_{K})$, and hence the search space can be reduced to $\mathcal{A}_{K}$. Finally, instead of computing a convex combination of $s^{t}$ and $w^{K,t-1}$ as in the classical Frank-Wolfe algorithm, the method optimizes over the convex hull of the new active set, namely
$\mathcal{S}^{t}$ $=\mathcal{S}^{t-1}\cup\{s^{t}\}$
$w^{t}$ $=\operatorname*{arg\,min}_{w\in\operatorname{conv}\mathcal{S}^{t}}h(w)$ (36)
Depending on the structure of the objective function, step (36) can be intractable or too costly to compute, which limits the use of FCFW in some practical applications. However, since $h$ is a quadratic here, (36) can be rewritten as
$\operatorname*{arg\,min}_{\beta\in\Delta_{t}}$ $\frac{1}{2}\left\lVert\sum_{l=0}^{t}\beta_{l}s^{l}-\frac{w^{K}}{n}\right\rVert^{2}.$ (37)
This a quadratic programming (QP) problem, which can be solved using classical off-the-shelf QP solvers osqp ; aps2019mosek ; ferreau2014qpoases ; bambade2023proxqp , or via first-order methods like gradient descent, since projection on the simplex can be efficiently computed. We summarize the full method in Algorithm 3.
Algorithm 3 Approximate Carathéodory via fully corrective Frank Wolfe
1:   Input: $w^{K}/n$, set $\mathcal{A}_{K}$, number of iterations $T$.
2:  Pick some $w\in\mathcal{A}_{K}$ and set $w^{K,0}=w$, $s^{0}=w^{K,0}$ and $\mathcal{S}^{0}=\{s^{0}\}$.
3:  for $t=1,\dots,T$ do
4:     Find $s^{t}$ such that
$s^{t}\in\operatorname*{arg\,min}_{s\in\mathcal{A}_{K}}s^{\top}(w^{K,t-1}-w^{K}/n).$
5:     Set $\mathcal{S}^{t}=\mathcal{S}^{t-1}\cup\{s^{t}\}$.
6:     Build $S_{t}\in\mathbb{R}^{p\times(t+1)}$, the matrix whose columns are the vectors in $\mathcal{S}^{t}$.
7:     Solve the following QP:
$\beta^{t}\in\operatorname*{arg\,min}_{\beta\geq 0,\mathbf{1}^{\top}\beta=1}$ $\frac{1}{2}\beta^{\top}S_{t}^{\top}S_{t}\beta-\left(\frac{w^{K}}{n}\right)^{\top}S_{t}\beta$
8:     $w^{K,t}=\sum_{l=0}^{t}\beta^{t}_{l}s^{l}$.
9:  end for
10:  return  $w^{K,T}$
We now give convergence guarantees on the distance between $w^{K,T}$ and $w^{K}/n$ for Algorithm 3. Linear convergence rates for conditional gradient methods usually depend either on the distance from the minimizer to the relative boundary of the feasible set guelat1986some ; garber2015faster , or on constants quantifying the conditioning of the feasible set, such as the pyramidal width lacoste2015global , the vertex-facet distance beck2017linearly , or the restricted width pena2016neumann . Those constants are notoriously difficult to apprehend, let alone compute in practical settings.
We give two convergence results, one based on the distance $r$ between $w^{K}/n$ and the relative boundary of $\operatorname{conv}\mathcal{A}_{K}$ garber2015faster , and the second one based on the notion of facial distance of the feasible set pena2019polytope , which turns out to be equivalent to the pyramidal width defined in lacoste2015global .
Definition 1
The facial distance of $\mathcal{A}_{K}$ is defined as
$\omega_{\mathcal{A}_{K}}:=\min_{\begin{subarray}{c}F\in\text{faces}(\operatorname{conv}\mathcal{A}_{K}),\\ \emptyset\subsetneq F\subsetneq\operatorname{conv}\mathcal{A}_{K}\end{subarray}}\text{dist}(F,\operatorname{conv}(\mathcal{A}_{K}\setminus F))$ (38)
We summarize the two convergence results in the next theorem.
Proposition 4.1
Let $w^{K,T}$ be the output of Algorithm 3 after $T$ iterations. Then
$\left\lVert w^{K,T}-\frac{w^{K}}{n}\right\rVert^{2}\leq\left\lVert w^{K,0}-\frac{w^{K}}{n}\right\rVert^{2}\left(1-\frac{\delta(\mathcal{A}_{K},w^{*})}{4D_{\mathcal{A}_{K}}^{2}}\right)^{T},$ (39)
where $D_{\mathcal{A}_{K}}$ is the diameter of $\operatorname{conv}(\mathcal{A}_{K})$ and
$\delta(\mathcal{A}_{K},w^{*}):=\max\left(\omega_{\mathcal{A}_{K}}^{2},(r^{*})^{2}\right),$
where $r^{*}$ is the radius of the largest relative ball centered at $\frac{w^{K}}{n}$ contained in $\operatorname{conv}(\mathcal{A}_{K})$, and $\omega_{\mathcal{A}_{K}}$ is the facial distance of $\mathcal{A}_{K}$ defined in (38).
Proof
Observe that the objective function $h$ is $1$-strongly convex and its gradient is $1$-Lipschitz, and its optimal value is 0. The rate with respect to the facial distance is then a consequence of (lacoste2015global, , Theorem 1) and the fact that the facial distance is equivalent to the pyramidal width (pena2019polytope, , Theorem 2). The rate with respect to the distance to the relative boundary is a consequence of (garber2015faster, , Section 4.2), and was also derived in combettes2023revisiting .
4.1.   Application to trimming the solution of the first stage
The output $w^{K,T}$ of Algorithm 3 is a convex combination of vectors in $\mathcal{A}_{K}$, i.e. we can write it as
$w^{K,T}=\sum_{l=0}^{T}\beta_{l}\begin{pmatrix}f_{i_{l}}(y_{i_{l}}^{k_{l}})\\ A_{i_{l}}y_{i_{l}}^{k_{l}}\\ e_{i_{l}}\end{pmatrix}=\sum_{i=1}^{n}\sum_{l\in I_{i}}\beta_{l}\begin{pmatrix}f_{i}(y_{i}^{k_{l}})\\ A_{i}y_{i}^{k_{l}}\\ e_{i}\end{pmatrix},$ (40)
where, as in (20), we regrouped the terms in $i$ by defining $I_{i}:=\{l\mid i_{l}=i\}$ for each $i=1,\dots,n$. For each $l$, define $\alpha_{l}=\frac{\beta_{l}}{\sum_{l^{\prime}\in I_{i_{l}}}\beta_{l^{\prime}}}$ and set
$\bar{w}^{K,T}=\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}\begin{pmatrix}f_{i}(y_{i}^{k_{l}})\\ A_{i}y_{i}^{k_{l}}\\ e_{i}\end{pmatrix}.$ (41)
The next theorem gives a bound on the quality of the approximation of $w^{K}$ by $\bar{w}^{K,T}$, and provides conditions under which $\bar{w}^{K,T}$ is well-defined. For clarity, and because it will be useful later on, we separate the first coordinate (the one corresponding to function values) from the coordinates corresponding to matrix vector multiplication.
{restatable}
lemApproxCarathSpecified Suppose we run Algorithm 1 for $T\geq n-1$ iterations with input $\mathcal{A}_{K}$ and vector $w^{K}/n$, which outputs $w^{K,T}$ defined in (40), and that we build $\bar{w}^{K,T}$ as in (41). Let
$h_{0}=\left\lVert w^{K,0}-\frac{w^{K}}{n}\right\rVert,$
where $w^{K,0}$ is the starting point of Algorithm 3. Let $D_{\mathcal{A}_{K}}$ be the diameter of $\mathcal{A}_{K}$ and let $\delta(\mathcal{A}_{K},\frac{w^{K}}{n})$ be as defined in Proposition 4.1. Assume that $T$ is large enough so that
$\frac{1}{n}>h_{0}\exp\left(-\frac{T}{8}\left(\frac{\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{D_{\mathcal{A}_{K}}^{2}}\right)\right).$ (42)
It then holds that $|I_{i}|\geq 1$ and that $\sum_{l\in I_{i}}\alpha_{l}=1$ for all $i$. Importantly here, $|I_{i}|>1$ for at most $T-n+1$ indices $i$. Moreover,
$\left\lvert\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})-w^{K}_{1}\right\rvert$ $\leq n\left(\sqrt{n}\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert+1\right)h_{0}\exp\left(-\frac{T\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{8D_{\mathcal{A}_{K}}^{2}}\right),$
$\left\lVert\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}A_{i}y_{i}^{k_{l}}-w^{K}_{2:m+1}\right\rVert$ $\leq n\left(\sqrt{n}\max_{i,y\in X_{i}}\left\lVert A_{i}y\right\rVert+1\right)h_{0}\exp\left(-\frac{T\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{8D_{\mathcal{A}_{K}}^{2}}\right).$
Proof
We start by showing that $|I_{i}|\geq 1$ for all $i$, which implies that $\alpha_{l}$ is well defined for all $l$. Indeed, the bound of Proposition 4.1 can be applied in particular to the last $n$ coordinates of $w^{K,T}-\frac{w^{K}}{n}$. This gives, for all $i=1,\dots,n$,
$\left(\sum_{l\in I_{i}}\beta_{l}-\frac{1}{n}\right)^{2}\leq h_{0}^{2}\left(1-\frac{1}{4}\left(\frac{\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{D_{\mathcal{A}_{K}}^{2}}\right)\right)^{T}\leq h_{0}^{2}\exp\left(-\frac{T}{4}\left(\frac{\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{D_{\mathcal{A}_{K}}^{2}}\right)\right).$
If, for some $i$, $I_{i}$ is empty, this implies that
$\frac{1}{n}\leq h_{0}\exp\left(-\frac{T}{8}\left(\frac{\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{D_{\mathcal{A}_{K}}^{2}}\right)\right),$
which contradicts our assumption. Therefore $|I_{i}|\geq 1$ for all $i=1,\dots,n$. Moreover, the definition of $w^{K,T}$ tells us that there are only $T+1$ terms in the sum, which implies that
$T+1=\sum_{i=1}^{n}|I_{i}|.$
Since $|I_{i}|\geq 1$ for all $i$, this implies that at $|I_{i}|>1$ for at most $T+1-n$ indices.
We now prove the inequalities. We start with the first one. The second one follows with the same argument. By the triangle inequality,
$\left\lvert\bar{w}^{K,T}-w^{K}_{1}\right\rvert$ $\leq\left\lvert\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{i_{l}})-nw^{K,T}_{1}\right\rvert+n\left\lvert w^{K,T}_{1}-\frac{w^{K}_{1}}{n}\right\rvert$
$=\left\lvert\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})-n\sum_{i=1}^{n}\sum_{l\in I_{i}}\beta_{l}f_{i}(y_{i}^{k_{l}})\right\rvert+n\left\lvert w^{K,T}_{1}-\frac{w^{K}_{1}}{n}\right\rvert$
The second term in the sum can be bounded using the rate from Proposition 4.1, which gives
$n\left\lvert w^{K,T}_{1}-\frac{w^{K}_{1}}{n}\right\rvert\leq nh_{0}\exp\left(-\frac{T\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{8D_{\mathcal{A}_{K}}^{2}}\right).$
For the first term, we have
$\left\lvert\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})-n\sum_{i=1}^{n}\sum_{l\in I_{i}}\beta_{l}f_{i}(y_{i}^{k_{l}})\right\rvert$ $=\left\lvert\sum_{i=1}^{n}\left(\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})-\sum_{l\in I_{i}}n\beta_{l}f_{i}(y_{i}^{k_{l}})\right)\right\rvert$
$=\left\lvert\sum_{i=1}^{n}\sum_{l\in I_{i}}(\alpha_{l}-n\beta_{l})f_{i}(y_{i}^{k_{l}})\right\rvert$
$\leq\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert\sum_{i=1}^{n}\sum_{l\in I_{i}}\left\lvert\alpha_{l}-n\beta_{l}\right\rvert$
$=\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert\sum_{i=1}^{n}\sum_{l\in I_{i}}\left\lvert\beta_{l}\left(\frac{1}{\sum_{l^{\prime}\in I_{i}}\beta_{l^{\prime}}}-n\right)\right\rvert$
$=\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert\sum_{i=1}^{n}\sum_{l\in I_{i}}\beta_{l}\left\lvert\frac{1}{\sum_{l^{\prime}\in I_{i}}\beta_{l^{\prime}}}-n\right\rvert$
Now, since $\sum_{l\in I_{i}}\beta_{l}>0$ and the term in absolute value does not depend on $l\in I_{i}$,
$\left\lvert\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{i_{l}})-n\sum_{i=1}^{n}\sum_{l\in I_{i}}\beta_{l}f_{i}(y_{i}^{i_{l}})\right\rvert$ $\leq\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert\sum_{i=1}^{n}\left\lvert\sum_{l\in I_{i}}\beta_{l}\left(\frac{1}{\sum_{l^{\prime}\in I_{i}}\beta_{l^{\prime}}}-n\right)\right\rvert$
$=\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert\sum_{i=1}^{n}n\left\lvert\frac{1}{n}-\sum_{l\in I_{i}}\beta_{l}\right\rvert$
$\leq\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert n\sqrt{n}\sqrt{\sum_{i=1}^{n}\left(\frac{1}{n}-\sum_{l\in I_{i}}\beta_{l}\right)^{2}}$
$\leq n\sqrt{n}\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert h_{0}\exp\left(-\frac{T\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{8D_{\mathcal{A}_{K}}^{2}}\right).$
The second to last inequality comes from the fact that $\left\lVert x\right\rVert_{1}\leq\sqrt{n}\left\lVert x\right\rVert_{2}$ for any $x\in\mathbb{R}^{n}$, while the last one comes from the rate given in Proposition 4.1.
Remark 6
One needs $|I_{i}|\geq 1$ for all $i$ for $\bar{w}^{K,T}$ to be well-defined. Inequality (42) provides a sufficient condition for it to be true. It is however hard to check in practice, since both $\delta(\mathcal{A}_{K},\frac{w^{K}}{n})$ and $D_{\mathcal{A}_{K}}$ are usually not available. Simple algebraic manipulations show that a sufficient condition for it to hold is
$n\geq\frac{8D_{\mathcal{A}_{K}}}{\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}\ln(nh_{0}).$
Since both $D_{\mathcal{A}_{K}}$ and $\delta(\mathcal{A}_{K},\frac{w^{K}}{n})$ depend on $n$, we cannot be sure that this holds, even as $n$ grows. In the considered practical application, we observed that one had $|I_{i}|\geq 1$ for all $i$ for values of $T$ smaller than $n+m$.
4.2.   Final reconstruction
From the decomposition of $\bar{w}^{K,T}$ as defined in (41), we are able to build a final solution to the original problem, just as in Section 3.3. We can moreover quantify the quality of the resulting solution by using the bound on the approximation of $w^{K}$ by $\bar{w}^{K,T}$ in Lemma 4.1. We first state the analogous of Lemma 3.2 in the case of approximate Carathéodory representations.
In order to lighten notations, we define the error terms induced by trimming using Algorithm 3 as
$\text{Err}_{f}(n,T)$ $:=n\left(\sqrt{n}\max_{i,y\in X_{i}}\left\lvert f_{i}(y)\right\rvert+1\right)h_{0}\exp\left(-\frac{T\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{8D_{\mathcal{A}_{K}}^{2}}\right)$
$\text{Err}_{A}(n,T)$ $:=n\left(\sqrt{n}\max_{i,y\in X_{i}}\left\lVert A_{i}y\right\rVert+1\right)h_{0}\exp\left(-\frac{T\delta(\mathcal{A}_{K},\frac{w^{K}}{n})}{8D_{\mathcal{A}_{K}}^{2}}\right),$
where $h_{0}$, $\delta(\mathcal{A}_{K},\frac{w^{K}}{n})$ and $D_{\mathcal{A}_{K}}$ are defined in Lemma 4.1.
Lemma 4.1
Let $v^{*}$ be the optimal value of (D) and (CoP). Assume that we run Algorithm 2 for $K$ iterations, followed by Algorithm 3 for $T\geq n-1$ iterations to trim the number of elements, outputting some $w^{K,T}$. Finally assume that we build $\bar{w}^{K,T}$ as in (41). Then, under the conditions of Lemma 4.1,
$\bar{w}^{K,T}_{1}\leq v^{*}+\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}+\text{Err}_{f}(n,T),$
$\left\lVert\bar{w}^{K,T}_{2:m+1}-b\right\rVert_{+}\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}+\text{Err}_{A}(n,T).$
Proof
Observe that
$\bar{w}^{K,T}_{1}-v^{*}$ $=(\bar{w}^{K,T}_{1}-w^{K}_{1})+(w^{K}_{1}-v^{*}),$
$\left\lVert\bar{w}^{K,T}_{2:m+1}-b\right\rVert_{+}$ $\leq\left\lVert\bar{w}^{K,T}_{2:m+1}-w^{K}_{2:m+1}\right\rVert_{+}+\left\lVert w^{K}_{2:m+1}-b\right\rVert_{+}.$
Since $z^{K}$, namely the output of Algorithm 2, consists of the first $1+m$ coordinates of $w^{K}$, we have in particular $w^{K}_{1}=z^{K}_{1}$ and $w^{K}_{2:m+1}=z^{K}_{2:m+1}$. Applying Theorem 3.1 then gives the bound on the second term of both equation. We can then bound $\bar{w}^{K,T}_{1}-w^{K}_{1}$ and $\left\lVert\bar{w}^{K,T}_{2:m+1}-w^{K}_{2:m+1}\right\rVert_{+}$ using Lemma 4.1.
Based on the previous lemmas, and using the decomposition (41), the exact same final reconstruction schemes as in Section 3.3 can be used. The theoretical guarantees obtained on the reconstruction are very similar, with two important differences. First, the dependence of the duality gap bound on the nonconvexity of the function is no longer governed by $m+1$, but rather by $T+1-n$, where $T$ is the number of iterations of the conditional gradient method Algorithm 3. This parameter $T$ can be controlled, and hence one can reduce the impact of the nonconvexity by choosing $T\leq n+m$. This comes at the price of additive terms $\text{Err}_{f}(n,T)$ and $\text{Err}_{A}(n,T)$ in the bounds, which are exponentially decreasing with $T$. We explicitly state and prove the result for convex domains in Corollary 4.1, which is the analogue of Theorem 3.2. The exact same holds for the analogue of Theorem 3.3, Theorem 3.3.2.2, Theorem 3.1, and we skip them in the sake of conciseness.
First let us state the following assumption, which summarizes the algorithmic framework we are working with.
Assumption 4.1
Let $v^{*}$ be the optimal value of (D) and (CoP). We assume that we run Algorithm 2 for $K$ iterations, followed by Algorithm 3 for $T\geq n-1$ iterations to trim the number of elements, outputting some $w^{K,T}$. We then build $\bar{w}^{K,T}$ as in (41), and we assume that the condition (42) of Lemma 4.1 holds.
4.2.1.   Convex domains
Recall that for convex domains, a natural candidate solution, based on the decomposition of $\bar{w}^{K,T}$ given in (41), is $\bar{x}\in\mathbb{R}^{d}$ defined as
$\bar{x}_{i}=\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}},\quad i=1,\dots,n.$ (43)
Corollary 4.1 (Analogue of Theorem 3.2)
Suppose Assumption 4.1 holds and that $X_{i}$ is convex for all $i=1,\dots,n$. Then, $\bar{x}$, as defined in (43), satisfies
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})$ $\leq v^{*}+\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}+\text{Err}_{f}(n,T)+(T+1-n)\max_{i}\rho(f_{i}),$
$\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-b\right\rVert_{+}$ $\leq\frac{2D_{\mathcal{C}}}{\sqrt{K+1}}+\text{Err}_{A}(n,T).$
Proof
Observe that
$\begin{split}\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-v^{*}&=\left(\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-\bar{w}^{K,T}_{1}\right)+\left(\bar{w}^{K,T}_{1}-v^{*}\right)\\ \left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-b\right\rVert_{+}&\leq\left\lVert\sum_{i=1}^{n}A_{i}\bar{x}_{i}-\bar{w}^{K,T}_{2:m+1}\right\rVert_{+}+\left\lVert\bar{w}^{K,T}_{2:m+1}-b\right\rVert_{+}\end{split}$ (44)
We can apply Lemma 4.1 to bound the last term of each equation. Moreover,
$\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-\bar{w}^{K,T}_{1}$ $=\sum_{i=1}^{n}f_{i}(\bar{x}_{i})-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})$
$=\sum_{i,|I_{i}|>1}\left(f_{i}\left(\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}}\right)-\sum_{l\in I_{i}}\alpha_{l}f_{i}(y_{i}^{k_{l}})\right)$
$\leq\sum_{i,|I_{i}|>1}\rho(f_{i})$
$\leq(T+1-n)\max_{i}\rho(f_{i}),$
and
$\sum_{i=1}^{n}A_{i}\bar{x}_{i}-\bar{w}^{K,T}_{2:m+1}$ $=\sum_{i=1}^{n}A_{i}\sum_{l\in I_{i}}\alpha_{l}y_{i}^{k_{l}}-\sum_{i=1}^{n}\sum_{l\in I_{i}}\alpha_{l}A_{i}y_{i}^{k_{l}}=0.$
Remark 7
In this section we have explored how one can use the fully-corrective Frank-Wolfe algorithm (FCFW) to efficiently compute approximate Carathéodory representation of the output $w^{K}$ of the first stage. It is important to note that other linearly convergent variants of Frank-Wolfe may be used. In particular, away-steps lacoste2015global Frank-Wolfe and the min-norm point (MNP) algorithm wolfe1976finding are natural candidates. The away-steps Frank-Wolfe variant did not yield good numerical performance, but the MNP algorithm did, so we mention it here. Like FCFW, it is a conditional gradient method relying on the same linear minimization step (35). The main difference with FCFW is that instead of optimizing over a convex hull at each iteration (as in (36)), it optimizes over the affine hull of an active set. This is very interesting, as this comes down to solving a linear system at each iteration, instead of solving a QP for FCFW (as in (37)). The linear system can be kept in triangular form and efficiently updated, which makes iterations of MNP very efficient in practice (see wolfe1976finding for the full algorithm description and implementation details). Moreover, MNP also exhibits a linear rate of convergence lacoste2015global , only worse than the rate of FCFW by a factor 2 in the exponent. More precisely, Proposition 4.1 still holds with
$\left\lVert w^{K,T}-\frac{w^{K}}{n}\right\rVert^{2}\leq\left\lVert w^{K,0}-\frac{w^{K}}{n}\right\rVert^{2}\left(1-\frac{\delta(\mathcal{A}_{K},w^{*})}{4D_{\mathcal{A}_{K}}^{2}}\right)^{T/2}.$ (45)
The remaining results of this section still hold, by simply replacing $T$ by $T/2$ in condition (42) and in the definitions of $\text{Err}_{f}(n,T)$ and $\text{Err}_{A}(n,T)$.
Finally, note that it might be possible to obtain approximate Carathéodory representations via other methods, such as simplicial decomposition ones holloway1974extension , of which the MNP algorithm is a subcase.
5.   Numerical results
In this section we explore the numerical performance of our method on two problems, a unit commitment problem and a problem of charging a fleet of electric vehicles. The code to reproduce the experiments is available at: https://github.com/bpauld/NonConvexOpt
5.1.   Unit Commitment problem
The unit commitment (UC) problem aims at finding a scheduling of power-generating units with minimum cost, subject to a demand constraint. We first describe the problem. For an in-depth survey of different UC problems and solvers, see padhy2004unit ; mallipeddi2014unit . We are given a network of $n$ power-generating units, and a time-horizon $N$. At each time step $t\in\{1,\dots,N\}$, unit $i$ can either be turned on or off, which is denoted by a binary variable $u^{i}_{t}\in\{0,1\}$. Turning on unit $i$ has cost $c_{01}^{i}$, and turning it off has cost $c_{10}^{i}$. The power output of unit $i$ at time step $t$ is written $g_{i}^{t}\in\mathbb{R}$. When $u^{i}_{t}=0$, unit $i$ is turned off and $g^{i}_{t}=0$. When $u^{i}_{t}=1$, the power generation $g^{i}_{t}$ lies between a minimum value $\underline{g}^{i}$ and a maximum value $\overline{g}^{i}$. As is standard for UC problems fattahi2017conic , we assume the cost of production to be a quadratic function of $g^{i}_{t}$. We assume that all units are turned off at time step 00. The cost functions can then be written
$f_{i}(u^{i},g^{i})$ $=\sum_{t=1}^{N}\ell_{i}(u^{i}_{t},g^{i}_{t}),\quad i=1,\dots,n,$ (46)
where
$\ell_{i}(u^{i}_{t},g^{i}_{t})=\begin{cases}c_{10}u^{i}_{t-1}&\text{ if }u^{i}_{t}=0,\\ c_{01}^{i}(1-u^{i}_{t-1})+\beta_{i}(g^{i}_{t})^{2}+\gamma_{i}g^{i}_{t}+\omega_{i}&\text{ if }u^{i}_{t}=1,\end{cases}$
where $u^{i}_{0}=0$ by assumption. The domain of the functions is
$\operatorname{dom}(f_{i})=\left\{(u^{i},g^{i})\in\{0,1\}^{N}\times\mathbb{R}^{N}\middle|\begin{array}[]{l}(u^{i}_{t},g^{i}_{t})=(0,0)\text{ or }\\ (u^{i}_{t},g^{i}_{t})\in(1,[\underline{g}^{i},\overline{g}^{i}]),\end{array}\text{ for all $t=1,\dots,N$}\right\}.$ (49)
Finally, the network must meet a demand $D_{t}\in\mathbb{R}$ at each time step $t$. The optimization of production costs can then be written as
$\begin{split}\mbox{minimize}&\ \sum_{i=1}^{n}f_{i}(u^{i},g^{i})\\ \text{subject to }&\ \sum_{i=1}^{n}g^{i}_{t}\geq D_{t},\quad t=1,\dots,N\\ &(u_{i},g_{i})\in\operatorname{dom}(f_{i}),\quad i=1,\dots,n.\end{split}$ (50)
This fits exactly withing the setting of (P) with $m=N$.
5.1.1.   Computing the conjugate
In this section we show how to compute the conjugate of the function $f_{i}$ defined in (46). For ease of notation, we drop all indices $i$ and superscripts $i$. Suppose we wish to compute $f^{*}(v,p)$ for $v\in\mathbb{R}^{N}$ and $p\in\mathbb{R}^{N}$. Let us define
$\Delta_{k,0}$ $=\max_{\begin{subarray}{c}u_{1},\dots,u_{k},\\ g_{1},\dots,g_{k},\\ u_{k}=0\end{subarray}}\left\{\sum_{t=1}^{k}\left(u_{t}v_{t}+g_{t}p_{t}\right)-\sum_{t=1}^{k}\ell(u_{t},g_{t})\right\},$
$\Delta_{k,1}$ $=\max_{\begin{subarray}{c}u_{1},\dots,u_{k},\\ g_{1},\dots,g_{k},\\ u_{k}=1\end{subarray}}\left\{\sum_{t=1}^{k}\left(u_{t}v_{t}+g_{t}p_{t}\right)-\sum_{t=1}^{k}\ell(u_{t},g_{t})\right\}.$
One then notices that $f^{*}(v,p)=\max\left(\Delta_{N,0},\Delta_{N,1}\right)$. Moreover, we have
$\Delta_{1,0}$ $=0,\quad\Delta_{1,1}=v_{1}-c_{01}+\max_{\underline{g}\leq g\leq\overline{g}}\left\{p_{1}g-\beta g^{2}-\gamma g-\omega\right\},$
and for $k>1$,
$\Delta_{k,0}$ $=\max\left(\Delta_{k-1,0},\Delta_{k-1,1}-c_{10}\right),$
$\Delta_{k,1}$ $=\max\left(\Delta_{k-1,0}-c_{01},\Delta_{k-1,1}\right)+v_{k}+\max_{\underline{g}\leq g\leq\overline{g}}\left\{p_{k}g-\beta g^{2}-\gamma g-\omega\right\}.$
The objective in the maximization steps is a one-dimensional quadratic over a bounded closed interval, so obtaining the optimal value is straightforward. Thus the value $f^{*}(v,p)$ can be obtained recursively, and the complexity is $O(N)$. Moreover, the gradient can be obtained by keeping track of the arguments yielding the maximum in the recursion.
5.1.2.   Linear minimization over the domain
When $\alpha_{k}=0$ in (16), one needs to compute a linear minimization over $\operatorname{dom}f_{i}$. We show how to do so now. As in the computation of the conjugate above, we drop all indices $i$. Given $(v,p)\in\mathbb{R}^{N}\times\mathbb{R}^{N}$, one must solve
$\min_{(u,g)\in\operatorname{dom}f}\sum_{t=1}^{N}(u_{t}v_{t}+g_{t}p_{t}).$
The structure of $\operatorname{dom}f$ tells us that the problem is separable, and the subproblems can be rewritten as
$\min_{u_{t},g_{t}}\quad$ $u_{t}v_{t}+g_{t}p_{t}$
subject to $(u_{t},g_{t})=(0,0)\text{ or }(u_{t},g_{t})\in(1,[\underline{g},\overline{g}]).$
Since the objective is linear, the minimum is necessarily attained at an extreme point, i.e. at one of $(0,0)$, $(1,\underline{g})$ or $(1,\overline{g})$.
5.1.3.   Implementation details
We generate random instances of the UC problem. Details on the data generation process can be found in Appendix B. To compute the optimal dual value $v^{*}$, we solve the dual (D) using projected gradient descent with a decreasing stepsize, which we stop when no more significant progress on the dual objective value is achieved. We run Algorithm 2 for $K=10^{4}$ iterations. We then either compute an exact Carathéodory representation (Algorithm 1), an approximate Carathéodory representation with FCFW (Algorithm 3), or an approximate Carathéodory representation with MNP (see Remark 7 and wolfe1976finding ). For Algorithm 3, we solve the QP using ProxQP bambade2023proxqp .
As we discussed in Remark 1, Assumption 2.4.2 holds for the unit commitment problem, namely for a point in the convex hull of the domain there always exist a a point in the domain which increases the overall electricity output. We write this solution $x^{F}$.
Finally, the bounds provided in Section 3.3 give solutions that are feasible up to an arbitrarily small term. Early experiments showed that solving a slightly perturbed version of the problem yielded fully feasible primal solutions. In the spirit of the approach described in paragraph 3.3.2.3, we thus apply our method to problem (50) perturbed by a vector $\theta\in\mathbb{R}^{N}$. Although the discussion in 3.3.2.3 suggests setting $\theta$ to $(N+1)\max_{i}\overline{g}^{i}$, this turned out to be unnecessarily large in our experiments. Instead, we set $\theta=\zeta\max_{i}\overline{g}^{i}$, with initial value $\zeta=1$. If the solution satisfies the demand, we stop there. Otherwise, we increase $\zeta$ by 1 and solve the new perturbed problem. We repeat this until the obtained solution meets the demand.
5.1.4.   Results
Table 1 presents the results for 10 randomly generated instances with $n=50$ units over $N=10$ time steps. While the theoretical duality gap is $(N+1)\max_{i}\gamma(f_{i})$, all of our experiments yield solutions whose duality gap is smaller than $\max_{i}\gamma(f_{i})$, a considerable improvement. Finally, the quality of the solutions is not significantly impacted by the Carathéodory approach (exact or approximate).
Table 1: Results for different random instances of the unit commitment problem with $n=50$ and $m=10$. $x^{F}$ corresponds to the reconstructed solution whose feasibility is better than the obtained convex combination, as in Equation 27. Exact Carathéodory corresponds to the case where the trimming is performed using Algorithm 1, while approximate Carathédodory representations are computed using either FCFW (Algorithm 3) or MNP (as mentionned in Remark 7). We see that the recovered solutions have duality gap much below the theoretical value of $(m+1)\max_{i}\gamma(f_{i})$.
Exact Carath. App. Carath. FCFW App. Carath. MNP
$v^{*}$ $\max_{i}\gamma(f_{i})$ $\zeta$ $f(x^{F})$ $\zeta$ $f(x^{F})$ $\zeta$ $f(x^{F})$
87 772 28 395 1 96 509 2 99 924 1 95 798
62 412 23 881 2 78 882 2 74 683 2 78 588
62 259 25 079 1 70 102 2 74 318 1 71 302
85 452 23 231 1 95 037 3 106 412 1 94 137
113 128 25 301 1 123 040 2 130 858 1 123 044
90 860 23 802 2 109 873 3 119 599 2 109 673
78 967 21 190 1 87 239 3 98 516 1 88 661
90 168 23 364 1 98 793 2 105 536 1 98 805
108 709 27 126 1 118 484 2 129 752 1 119 376
44 832 22 921 1 52 141 2 53 397 1 51 271
We now analyze the running time of our method. To do so, we generate random instances of the UC problem with $N=20$ and different values of $n$. We then run Algorithm 2 for $K=10^{4}$.
Figure 1 (left) plots the average runtime of Algorithm 2 (over 5 runs). We see that there is a linear relationship between the number of units $n$ and the running time. This is expected, as the bottleneck of Algorithm 2 is the computation of the linear minimization oracle at each point $i\in\{1,\dots,n\}$. Although the running time may seem high at first, there are important things to consider here. First, the computation of the conjugate described in Section 5.1.1 is an iterative approach, implemented with for loops in our code, which are notoriously slow in Python. One could thus highly benefit from an implementation in another programming language. Moreover, we implemented our method on a single machine but the structure of the linear minimization step (15) make it fully parallelizable, so one could further take advantage of having several machines solving (16) in parallel.
Figure 1 (middle and right) plots the average running time of the different Carathéodory algorithms. We see on the middle plot that the running time of the exact algorithm (Algorithm 1) becomes prohibitively large even for moderate $n$, and hence we did not run it for values over $n=200$. The right plot shows that the running time of the MNP algorithm is better than the one of FCFW. This is expected, as each iteration of of FCFW solves a QP whose size increases with the number of iterations, and one must run at least $n$ iterations in order to construct a final solution. This plot and the result from Table 1 lead us to conclude that the best method to obtain Carathéodory representations is the MNP algorithm.
Figure 1: Running time of the different algorithms with respect to $n$, averaged over 5 runs. (Left) Running time of Algorithm 2. (Middle) Running time of all three Carathéodory approaches. We see that the exact approach (Algorithm 1) has long running time even for small values of $n$. (Right) Running time of the approximate Carathéodory approaches for larger values of $n$. We see that as $n$ grows, the MNP algorithm becomes more efficient than the FCFW algorithm.
5.2.   Charging of Plug-in Electric Vehicles
Inspired by vujanic2016decomposition , we now consider the problem of efficiently charging a fleet of Plug-in Electric Vehicles (PEVs). Let us describe the problem setting. We are given a fleet of $n$ vehicles, which must be charged over $N$ intervals of time, each of length $\Delta$. The charge level of vehicle $i$ at time step $k$ is denoted as $e_{i}^{k}$. Whether or not vehicle $i$ is charging at time step $k$ is given by a variable $u^{i}_{k}\in\{0,1\}$. The charge level at time step $k+1$ is then given by
$e^{i}_{k+1}=e^{i}_{k}+P_{i}\Delta\xi_{i}u^{i}_{k},$
where $P_{i}$ is the charging rate at station $i$ and $\xi_{i}$ is the charging conversion efficiency. The initial charge level is given by $E_{\text{init}}^{i}$, and the final charge must be at least $E_{\text{ref}}^{i}$ for all $i$. Moreover, the level of charging of vehicle $i$ cannot exceed a maximum value $E_{\text{max}}^{i}$. Finally, the system cannot impose too much stress on the overall network. This translates into the constraint
$\sum_{i=1}^{n}P_{i}u^{i}_{k}\leq P^{\text{max}}_{k},\ k=1,\dots,N,$
for some given $P^{\text{max}}\in\mathbb{R}^{N}$. If $C^{u}\in\mathbb{R}^{N}$ denotes the price vector for electricity consumption at each time step, we can write the problem of optimizing costs as
minimize $\ \sum_{i=1}^{n}f_{i}(u^{i})$
subject to $\ \sum_{i=1}^{n}P_{i}u^{i}_{k}\leq P^{\text{max}}_{k},\quad k=1,\dots,N,$
$u^{i}\in\operatorname{dom}(f_{i}),\quad i=1,\dots,n,$
where
$f_{i}(u^{i})=\sum_{k=1}^{N}P_{i}C^{u}_{k}u^{i}_{k}$
and
$\operatorname{dom}f_{i}=\left\{u\in\{0,1\}^{N}\ \middle|\begin{array}[]{l}E_{\text{init}}^{i}+\sum_{k=1}^{N}P_{i}\Delta\xi_{i}u_{k}\geq E_{\text{ref}}^{i},\\ E_{\text{init}}^{i}+\sum_{k=1}^{j}P_{i}\Delta\xi_{i}u_{k}\leq E_{\text{max}}^{i},\ j=1,\dots,N\end{array}\right\}$
The computation of the Fenchel conjugate of $f_{i}$ simply consists of a greedy strategy vujanic2016decomposition , so we omit it here.
5.2.1.   Data generation
We set $n=500$ and $N=24$ and refer the reader to (vujanic2016decomposition, , Appendix B) for the full data generation procedure.
5.2.2.   Implementation details
We implement our method with the MNP algorithm to trim the number of elements in the output of Algorithm 2. We compare it with the randomized algorithm from udell2016bounding , the only other method we are aware of achieving similar duality gap bounds as our method. Since the domain of the function is not convex, we resort to the perturbation approach described in 3.3.2.3 and hence apply our approach to the perturbed problem ($\text{CoP}_{\theta}$). Setting $\theta=N\max_{i}P_{i}$ ensures that the assumptions of 3.3.2.3 are satisfied vujanic2016decomposition . We build the final solution as in (30). We also do this for the randomized algorithm from udell2016bounding .
5.2.3.   Results
We explore how the feasibility and the function values evolve as a function of $K$, the number of iterations of Algorithm 2. We write $\bar{x}_{K}$ as the final reconstructed solution.
Figure 2 plots the average unfeasibility (over 10 runs) of $\bar{x}_{K}$ in the perturbed problem ($\text{CoP}_{\theta}$) (left) and in the original primal problem (P) (right). First we see that for the original primal problem, a small value of $K=10^{3}$ was sufficient for all solutions to be fully primal feasible. The left plot shows that infeasibility in the perturbed primal problem behaves similarly to the (scaled) $O(1/\sqrt{K})$ curve in red, as is expected by our theoretical bounds. We hypothesize that the slightly worse than $O(1/\sqrt{K})$ behavior is due to the additive exponential error term $\text{Err}_{A}(n,T)$ which appears in the bounds when using approximate Carathéodory algorithms (c.f. Lemma 4.1).
Figure 2: Infeasibility of the final reconstructed solution $\bar{x}_{K}$ with respect to the number of iterations $K$ of Algorithm 2, averaged over 10 runs. (Left) Infeasibility of the solution in the perturbed primal problem. In red is just the (scaled) curve $1/\sqrt{K}$ for reference. We see that the infeasibility seems to be decreasing close to the $O(1/\sqrt{K})$ rate as predicted by our bounds, and the slightly worse behavior is most likely due to the additive exponential error term $\text{Err}_{A}(n,T)$. (Right) Infeasibility of the solution in the original primal problem. We see that over 10 runs, it was always sufficient to run for only $K=1000$ iterations to obtain a fully feasible solution.
Figure 3 plots the evolution of the function values for two specific instances of the problem. In orange is the function value for the output of the randomized algorithm from udell2016bounding , while the curve in blue is the output of our method for different $K$. Finally, the vertical dashed red line represents the first value of $K$ for which the original primal problem is fully primal feasible for $\bar{x}_{K}$. At first sight, it might seem strange that the value of the function increases with $K$, but this is actually expected from the left plot in Figure 2. Indeed, for small values of $K$, the perturbed primal problem (which is the one we numerically solve) is more infeasible than for large values of $K$, and thus smaller objective function values can be obtained. One is thus tempted to take small values of $K$, but not too small so that the original primal problem is fully feasible. In that sense, the ‘optimal’ value of $K$ is the one given by the red vertical line, i.e. the smallest value of $K$ for which the original primal problem is fully feasible.
For large-scale problems, we saw in the experiments on the unit commitment problem that the main bottleneck is the running time of the first algorithm Algorithm 2. This suggests the following strategy for the method. While running Algorithm 2, one should periodically stop the algorithm at some iteration $K$ with iterate $z_{K}$, compute the final solution $\bar{x}_{K}$ using the approximate Carathéodory algorithm, and check if that solution is feasible in the original primal problem. If it is, then one can stop the whole procedure. Otherwise, one can go back to Algorithm 2 and resume from $z_{K}$, and do another check later on. Since the running time of approximate Carathéodory algorithms is much lower than Algorithm 2, this will not significantly impact the running time of the overall procedure.
Figure 3: Typical behavior of our method on two specific instances of the PEVs problem. The curve in blue represents the value of the function at the final reconstructed solution $\bar{x}_{K}$ with respect to the number of iterations $K$ of Algorithm 2. In orange is the value obtained by the randomized algorithm from udell2016bounding . The vertical dashed red line indicates the first value of $K$ for which $\bar{x}_{K}$ is primal feasible, i.e. $A\bar{x}_{K}-b\leq 0$.
Appendix A Technical lemmas
[…]
To prove the above lemma, we prove the following, stronger result. Recall that a function $f$ is 1-coercive if
$\lim_{\left\lVert x\right\rVert\rightarrow\infty}\frac{f(x)}{\left\lVert x\right\rVert}=\infty.$
In particular, a function with a compact domain is 1-coercive.
Lemma A.1
Suppose $f$ is proper, closed and 1-coercive. Then
$\left\{(Ax,f^{**}(x))\mid x\in\operatorname{dom}f^{**}\right\}\subset\operatorname{conv}\left\{(Ax,f(x))\mid x\in\operatorname{dom}f\right\}.$
Moreover, if for some $\tilde{x}$,
$(A\tilde{x},f^{**}(\tilde{x}))\in\operatorname{ext}\left\{(Ax,f^{**}(x))\mid x\in\operatorname{dom}f^{**}\right\},$
then $\tilde{x}\in\operatorname{dom}f$ and $f^{**}(\tilde{x})=f(\tilde{x})$.
Proof
Let $z=(A\tilde{x},f^{**}(\tilde{x}))\in\left\{(Ax,f^{**}(x))\mid x\in\operatorname{dom}f^{**}\right\}$.
We have $(\tilde{x},f^{**}(\tilde{x}))\in\operatorname{epi}f^{**}=\operatorname{conv}\operatorname{epi}f$ (hiriart1996convex, , Theorem X.1.5.3). Therefore, for some $K\in\mathbb{N}$, $\alpha\in\Delta_{K}$ and some $x_{j}\in\operatorname{dom}f$,
$(\tilde{x},f^{**}(\tilde{x}))=\sum_{j=1}^{K}\alpha_{i}(x_{j},t_{j})\ \text{ with }\ f(x_{j})\leq t_{j}$ (51)
and thus
$z=(A\tilde{x},f^{**}(\tilde{x}))=\sum_{j=1}^{K}\alpha_{j}(Ax_{j},t_{j})\ \text{ with }\ f(x_{j})\leq t_{j}.$
In particular, we have
$f^{**}(\tilde{x})=f^{**}(\sum_{j=1}^{K}\alpha_{j}x_{j})\leq\sum_{j=1}^{K}\alpha_{j}f^{**}(x_{j})\leq\sum_{j=1}^{K}\alpha_{j}f(x_{j})\leq\sum_{j=1}^{K}\alpha_{j}t_{j}=f^{**}(\tilde{x}).$ (52)
All the above inequalities are therefore equalities, and in particular,
$(A\tilde{x},f^{**}(\tilde{x}))=\sum_{j=1}^{K}\alpha_{j}(Ax_{j},f(x_{j}))$ (53)
so that $z\in\operatorname{conv}\left\{(Ax,f(x))\mid x\in\operatorname{dom}f\right\}$ which proves the first statement. Up to regrouping, we can assume that $(Ax_{j_{1}},f^{**}(x_{j_{1}}))\not=(Ax_{j_{2}},f^{**}(x_{j_{2}}))$ for all $j_{1}\not=j_{2}$ in the sum.
Now suppose that $z$ is an extreme point of the set $\left\{(Ax,f^{**}(x))\mid x\in\operatorname{dom}f^{**}\right\}$. Recall that $z=(A\tilde{x},f^{**}(\tilde{x}))$ and observe that (51) and (52) tell us that
$z=(A\tilde{x},f^{**}(\tilde{x}))=\sum_{j=1}^{K}\alpha_{j}(Ax_{j},f^{**}(x_{j})).$
Since $z$ is an extreme point, the above convex combination must be trivial, i.e. we must have $\alpha_{j}=0$ for all but one index. Suppose without loss of generality that $\alpha_{1}=1$ and that $\alpha_{j}=0$ for $j>1$. We then have, by (51), $\tilde{x}=x_{1}\in\operatorname{dom}f$. We also have, by (53), $f^{**}(\tilde{x})=f(x_{1})=f(\tilde{x})$.
Lemma A.2
Suppose $f$ is proper, closed, and 1-coercive. Then
$\left\{\begin{pmatrix}f^{**}(x)\\ Ax\end{pmatrix}\mid x\in\operatorname{dom}f^{**}\right\}+\mathbb{R}^{m+1}_{+}=\operatorname{conv}\left\{\begin{pmatrix}f(x)\\ Ax\end{pmatrix}\mid x\in\operatorname{dom}f\right\}+\mathbb{R}^{m+1}_{+}$
Proof
Let $z\in\left\{\begin{pmatrix}f^{**}(x)\\ Ax\end{pmatrix}\mid x\in\operatorname{dom}f^{**}\right\}+\mathbb{R}^{m+1}_{+}$ so that
$z=\begin{pmatrix}f^{**}(\tilde{x})\\ A\tilde{x}\end{pmatrix}+\begin{pmatrix}\delta_{1}\\ \delta_{2}\end{pmatrix},$
with $\delta_{1}\in\mathbb{R}_{+},\delta_{2}\in\mathbb{R}_{m}^{+}$. It is then clear that
$\begin{pmatrix}\tilde{x}\\ z\end{pmatrix}\in\left\{(x,r_{0},r)\mid f^{**}(x)\leq r_{0},Ax\leq r\right\}.$
Using (lemarechal2001geometric, , Corollary A.6), this implies actually that
$\begin{pmatrix}\tilde{x}\\ z\end{pmatrix}\in$ $\operatorname{cl}\operatorname{conv}\left\{(x,r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\right\}$
$=\operatorname{conv}\left\{(x,r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\right\}$
where the equality is due to Lemma A.3. In particular, we have
$z$ $\in\operatorname{conv}\left\{(r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\text{ for some }x\right\}$
$=\operatorname{conv}\left\{\left\{\begin{pmatrix}f(x)\\ Ax\end{pmatrix}\mid x\in\operatorname{dom}f\right\}+\mathbb{R}^{m+1}_{+}\right\}$
$=\operatorname{conv}\left\{\begin{pmatrix}f(x)\\ Ax\end{pmatrix}\mid x\in\operatorname{dom}f\right\}+\mathbb{R}^{m+1}_{+}.$
Now suppose
$z=\begin{pmatrix}z_{1}\\ z_{2}\end{pmatrix}$ $\in\operatorname{conv}\left\{\begin{pmatrix}f(x)\\ Ax\end{pmatrix}\mid x\in\operatorname{dom}f\right\}+\mathbb{R}^{m+1}_{+}$
$=\operatorname{conv}\left\{(r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\text{ for some }x\right\}.$
Then there exists $\alpha\geq 0$ such that $11^{\top}\alpha=$ 1 and
$(z_{1},z_{2})=\sum_{k}\alpha_{k}(f(x_{k}),Ax_{k})+(\delta_{1},\delta_{2})$
for some $\delta_{1}\in\mathbb{R}_{+},\delta_{2}\in\mathbb{R}_{+}^{m}$ and some $x_{k}\in\operatorname{dom}f$. Let us set $y=\sum_{k}\alpha_{k}x_{k}$. Then
$\begin{pmatrix}y\\ z\end{pmatrix}$ $\in\operatorname{conv}\left\{(x,r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\right\}$
$=\operatorname{cl}\operatorname{conv}\left\{(x,r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\right\}$
=$=\left\{(x,r_{0},r\mid f^{**}(x)\leq r_{0},Ax\leq r\right\}$
where the first equality is by Lemma A.3 and the second is again from (lemarechal2001geometric, , Corollary A.6). This ends the proof.
Lemma A.3
Suppose $f$ is proper, closed, and 1-coercive. Then
$\operatorname{conv}\left\{(x,r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\right\}$
is closed.
Proof
Suppose $(x_{k},r_{0,k},r_{k})\in\operatorname{conv}\left\{(x,r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\right\}$ converges to $(\tilde{x},\tilde{r}_{0},\tilde{r})$. We want to show that $(\tilde{x},\tilde{r}_{0},\tilde{r})\in\operatorname{conv}\left\{(x,r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\right\}$. For all $k$, there exists $\alpha_{k}\geq 0$ such that $11^{\top}\alpha_{k}=$ 1 and
$(x_{k},r_{0,k},r_{k})=\sum_{j}\alpha_{kj}(x_{kj},r_{0,kj},r_{kj})$
with $(x_{kj},r_{0,kj})\in\operatorname{epi}f$ and $Ax_{kj}\leq r_{kj}$. In particular, this implies that $(x_{k},r_{0,k})\in\operatorname{conv}\operatorname{epi}f$ for all $k$. Since $\operatorname{conv}\operatorname{epi}f$ is closed (hiriart1996convex, , Lemma X.1.5.3), this implies that $(\tilde{x},\tilde{r}_{0})\in\operatorname{conv}\operatorname{epi}f$. Therefore there exists $(x_{p},r_{0,p})\in\operatorname{epi}f$ such that
$(\tilde{x},\tilde{r}_{0})=\sum_{p}\beta_{p}(x_{p},r_{0,p})$
with $\beta\geq 0$, $11^{\top}\beta=$ 1. Moreover, we have
$Ax_{k}=\sum_{k}\alpha_{kj}Ax_{kj}\leq\sum_{k}\alpha_{kj}r_{kj}=r_{k}.$
Taking limits, this implies that $A\tilde{x}\leq\tilde{r}$. Define $r_{p}=Ax_{p}+(\tilde{r}-A\tilde{x})$. Then $Ax_{p}\leq r_{p}$ and $\sum_{p}\beta_{p}r_{p}=A\sum_{p}\beta_{p}x_{p}+(\tilde{r}-A\tilde{x})=\tilde{r}$. Thus
$(\tilde{x},\tilde{r}_{0},\tilde{r})\in\operatorname{conv}\left\{(x,r_{0},r)\mid f(x)\leq r_{0},Ax\leq r\right\}$
Appendix B Experimental details
In this section, we give details on the data generation of Section 5. We write $U(c,d)$ as the uniform distribution in the interval $[c,d]$. At each time step $t$, the demand $D_{t}$ is generated according to $D_{t}\sim U(100,300)$. For each unit $i$, we then set the maximum and minimum generation parameters $\overline{g}^{i}$ and $\underline{g}^{i}$ as
$p^{i}$ $\sim\frac{1}{n}\times U(100,300)$
$\overline{g}^{i}$ $=2p^{i}$
$\underline{g}^{i}$ $=0.5p^{i}.$
We then set, for all units $i$,
$\beta_{i}$ $\sim U(1,20),\quad\gamma_{i}\sim U(3,5),\quad\omega_{i}\sim U(30,50),$
$c_{01}^{i}$ $=\frac{\sum_{j}\beta_{j}(p^{j})^{2}+\gamma_{j}p^{j}+\omega_{j}}{2n},\quad c_{10}^{i}=\frac{c_{01}^{i}}{4}.$
We found that choosing the parameters this way allows for interesting settings where the choice of when to turn on/off units of production, and of how much power should be generated, is non-trivial.
Funding
This work was funded in part by the French government under management of Agence Nationale de la recherche as part of the “Investissements d’avenir” program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute) and a Google focused award.
Conflict of interest
The authors declare that they have no conflict of interest.